### v1

# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

Abstract
We study the computation of lower and upper probabilities of hitting a target set of states for imprecise Markov chains. For these, transition uncertainty is modelled by a convex set of transition matrices. In the precise case, hitting probabilities are the minimal nonnegative solution of a linear system and admit a closed-form expression. We study the notion of reachability in the imprecise setting. The literature review highlights few different definitions of lower reachability; thus we explore the relations among them, presenting examples to clarify their logical implications. Using this revised definition of reachability for imprecise Markov chain, we partition the state space into classes of states whose hitting probabilities are trivially zero or one and those which require further computation. For these nontrivial states, we show that lower and upper hitting probabilities are the unique solutions of two nonlinear fixed-point equations. For the practical computation of lower and upper hitting probabilities, we propose iterative algorithms that alternate between solving a linear system and choosing an extreme point from the set of transition matrices. Numerical experiments demonstrate that, in practice, these algorithms converge in substantially fewer iterations than the theoretically established worst-case bound.
keywords:
Imprecise Markov Chain , Hitting Probability , Imprecise Probability
journal: International Journal of Approximate Reasoning
[…]
organization=Eindhoven University of Technology, city=Eindhoven, country=Netherlands
1 Introduction
Hitting probabilities play a central role in the analysis of stochastic systems, capturing the likelihood that a process enters a designated set of target states in some finite time. In the classical precise setting of homogeneous Markov chains, these probabilities arise as the minimal nonnegative solution of a linear system and admit a closed-form representation via the fundamental matrix [norrisbook1997markov, kemeny1976finite]. Hitting probabilities can also be regarded as discrete harmonic functions and enjoy a deep analogy with node voltages in potential theory and electrical networks [DoyleSnell1984].
In practical applications there is often uncertainty about the transition dynamics, which can be estimated from limited data or elicited from domain experts, yielding a convex set of transition matrices rather than a single, exact one. A natural and rigorous way to represent this kind of epistemic uncertainty is through credal sets – convex sets of probability distributions – leading to the framework of imprecise Markov Chains (IMCs) [DeCooman2009, HermansSkulj2014, hartfiel_seneta_1994]. In this framework, each row of the transition matrix is specified by a convex set of admissible probability distributions, and the inference quantities of interest are lower and upper hitting probabilities. These bound the true probability of reaching the target states under all compatible transition matrices. These values provide conservative guarantees that are crucial in high-stakes decision-making contexts. They naturally connect to robust control, since they provide worst-case bounds used in min–max robust MDPs under transition uncertainty [NilimElGhaoui2005], and to stochastic-game formulations, as they correspond to value bounds in zero-sum reachability games [Bovy2024].
Our contributions build on the work of Krak et al. [krak2019hitting, krak2021comphit]. They developed characterisations of hitting times and probabilities for IMCs under various model semantics and presented algorithms for computing lower and upper hitting times. We extend this work by (i) discussing and clarifying the definition of reachability in the imprecise setting, (ii) giving a more general characterisation of lower and upper hitting probabilities and (iii) proposing iterative algorithms to compute these quantities.
The structure of the paper is as follows. Section 2 introduces the preliminary notions needed for the following sections. In particular, we define hitting probabilities for homogeneous Markov chains, we explain the concept of imprecise Markov chains, and we introduce lower and upper transition operators. Section 3 begins our contributions. There, we formalise the notions of lower and upper reachability in the imprecise setting; we review the different definitions that appear in the literature, clarify the logical relationships between them, and give a few illustrative examples. With these definitions in hand, we isolate those states whose (lower/upper) hitting probabilities are trivial, either zero or one. We then prove that for nontrivial states, lower and upper hitting probabilities are the unique solutions of two nonlinear fixed-point equations.
In Section 4, we turn our attention to developing iterative algorithms to compute lower and upper hitting probabilities. Inspired by the work of Krak et al. [krak2021comphit] on hitting times, we propose iterative algorithms to compute lower and upper hitting probabilities that alternate between solving a linear system (restricted to a suitable subspace) and selecting an extreme point transition matrix from the credal sets. We rigorously prove their convergence and, in Section 5, we provide a detailed analysis of their computational cost. We conclude the paper with Section 6 which is dedicated to numerical experiments. Here we show how the theoretical bound on the number of iterations previously established is never attained in practice, and we point out an interesting relation between the density of the graph and the average number of iterations needed for convergence.
2 Preliminaries
This section introduces the basic concepts of stochastic processes [Dudley2002], Markov chains [norrisbook1997markov, kemeny1976finite], hitting probabilities for homogeneous Markov chains [norrisbook1997markov], and imprecise Markov chains [DeCooman2009, HermansSkulj2014, hartfiel_seneta_1994].
2.1 Stochastic Processes and Markov Chains
Let us denote by $\mathbb{N}$ the set of positive integers and define $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$. Let $\mathcal{X}$ be a discrete nonempty finite set of cardinality $N\in\mathbb{N}$ and denote by $2^{\mathcal{X}}$ its power set. Fix an arbitrary ordering over $\mathcal{X}$: this allows us to identify functions in $\mathbb{R}^{\mathcal{X}}$ and $N$-dimensional vectors.
Define the sample space $\Omega\coloneqq\mathcal{X}^{\mathbb{N}_{0}}$ and endow it with the product $\sigma$-algebra $\mathcal{F}\coloneqq(2^{\mathcal{X}})^{\otimes\mathbb{N}_{0}}$, i.e. the $\sigma$-algebra generated by cylinder sets. A stochastic process $(X_{n})_{n\in\mathbb{N}_{0}}$ is a family of random variables on this measurable space
$X_{n}:(\Omega,\mathcal{F})\to(\mathcal{X},2^{\mathcal{X}})$
where we set $X_{n}(\omega)=\omega(n)$ for all $\omega\in\Omega$ and all $n\in\mathbb{N}_{0}$ [Dudley2002]. Let $\mathbb{P}$ be a probability measure on the measurable space $(\Omega,\mathcal{F})$, so that $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The probability measure $\mathbb{P}$ determines the law of the stochastic process; in the sequel, we will identify the process with the probability measure $\mathbb{P}$, which allows us to consider different processes $\mathbb{P}$ over the same measurable space $(\Omega,\mathcal{F})$ [krak2019hitting].
A stochastic process $\mathbb{P}$ is said to be a Markov chain if it satisfies the Markov property:
$\mathbb{P}(X_{n+1}=x_{n+1}\mid X_{0:n}=x_{0:n})=\mathbb{P}(X_{n+1}=x_{n+1}\mid X_{n}=x_{n})$ (1)
for all $x_{0},\dots,x_{n},x_{n+1}\in\mathcal{X}$ and all $n\in\mathbb{N}_{0}$. This property is often interpreted as the process being memoryless; the transition probabilities depend only on the current state, and any previous history is not relevant. A Markov chain $\mathbb{P}$ is said to be (time-)homogeneous if
$\mathbb{P}(X_{n+1}=x\mid X_{n}=y)=\mathbb{P}(X_{1}=x\mid X_{0}=y)$ (2)
for all $x,y\in\mathcal{X}$ and for all $n\in\mathbb{N}_{0}$.
A matrix $T\in\mathbb{R}^{N\times N}$ is called a transition matrix if it satisfies
  1. 1.
    $T(x,y)\geq 0$ for all $x,y\in\mathcal{X}$, and
  2. 2.
    $\sum_{y\in X}T(x,y)=1$ for all $x\in\mathcal{X}$.
Note that each row $T(x,\cdot)$ of a transition matrix is a probability mass function on $\mathcal{X}$. Every transition matrix $T$ uniquely determines a homogeneous Markov chain $\mathbb{P}_{T}$ up to its initial distribution, by setting
$\mathbb{P}_{T}(X_{n+1}=x\mid X_{n}=y)=T{(x,y)}$ (3)
for all $x,y\in\mathcal{X}$ and for all $n\in\mathbb{N}_{0}$. Thus, transition matrices are often used as the canonical parameter to specify homogeneous Markov chains.
Definition 2.1 (Reachability [norrisbook1997markov]).
Let $\in\mathcal{X}, xyx\neq y$, and let $T$ be a transition matrix. We say that $x$ reaches $y$, and we denote it with $\to^{T}$”, if there exists $n\in\mathbb{N}$ and a sequence of states $x=x_{0},\ldots,x_{n}=y$ in $\mathcal{X}$ such that $T(x_{k},x_{k+1})>0$ for all $0\leq k\leq n-1$, and $x_{k}\neq x_{j}$ for all $0\leq k<j\leq n$.
The following proposition gives the equivalence between three formulations of reachability for homogeneous Markov chains.
Proposition 2.2 ([norrisbook1997markov, Theorem 1.2.1]).
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain defined on the state space $\mathcal{X}$ with transition matrix $T$. Let $x,y\in\mathcal{X}$. Then the following statements are equivalent:
  1. 1.
    $\mathbb{P}_{T}(X_{n}=y\mid X_{0}=x)>0$, for some $n\in\mathbb{N}_{0}$,
  2. 2.
    $x\to^{T}y$,
  3. 3.
    $T^{n}(x,y)>0$ for some $n\in\mathbb{N}_{0}$.
2.2 Hitting Probabilities for Markov Chains
In this section, we introduce hitting indicators and hitting probabilities, then we characterise hitting probabilities for homogeneous Markov chains.
Definition 2.3.
Let $A\subseteq\mathcal{X}$ be a nonempty target set of states. The hitting indicator is defined as the function $\phi_{A}:\Omega\to\{0,1\}$ given by
$\phi_{A}(\omega)\coloneqq\sup_{n\in\mathbb{N}_{0}}\,\mathbb{1}_{A}(\omega(n)),$
where $\mathbb{1}_{A}$ is the indicator function for $A$.
This quantity is a Boolean function having value one if the path reaches $A$ and zero otherwise.
Definition 2.4.
We define the hitting probability of a process $\mathbb{P}$ starting in $x\in\mathcal{X}$ as the expected value of the hitting indicator $\mathbb{E}_{\mathbb{P}}[\phi_{A}\mid X_{0}=x].$
The hitting probability of a process $\mathbb{P}$ starting in $x\in\mathcal{X}$ can also be seen as
$\mathbb{P}\,\Big(\exists n\in\mathbb{N}:X_{n}\in A\mid X_{0}=x\Big)=\mathbb{P}\,\Bigg(\bigcup_{n\in\mathbb{N}_{0}}\{X_{n}\in A\}\ \Big|\ X_{0}=x\Bigg),$
i.e. the probability that $A$ is reached in some number of steps. In the remainder of the paper, we collect these hitting probabilities in the real-valued function $p_{A}^{\mathbb{P}}:\mathcal{X}\to[0,1]$ defined by
$p_{A}^{\mathbb{P}}(x)\coloneqq\mathbb{E}_{\mathbb{P}}[\phi_{A}\mid X_{0}=x].$
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain with transition matrix $T$, and define $p^{T}\coloneqq p_{A}^{\mathbb{P}_{T}}$. Let us introduce the set of states that cannot reach $A$, namely
$\mathcal{C}_{T}\coloneqq$ $\{x\in A^{c}\mid\forall y\in A:x\not\to^{T}y\}$
$=$ $\{x\in A^{c}\mid\forall y\in A,\forall n\in\mathbb{N}_{0},:T^{n}(x,y)=0\}.$
We can characterise the hitting probabilities of a homogeneous Markov chain as the minimal nonnegative solution to a linear system of equations.
Proposition 2.5 ([norrisbook1997markov, Theorem 1.3.2]).
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain defined on the state space $\mathcal{X}$ with transition matrix $T$. Then $p^{T}(x)=1$ for all $x\in A$, and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. Moreover, $p^{T}$ is the (unique) minimal nonnegative solution of the system
$p^{T}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T\,p^{T}.$ (4)
2.3 Restrictions and Extensions of Functions and Matrices to a Subspace
Let $\emptyset\subset\mathcal{S}\subseteq\mathcal{Y}\subseteq\mathcal{X}$ and $f\in\mathbb{R}^{\mathcal{Y}}$. The definitions below give a canonical way to view functions and operators on the smaller set $\mathcal{S}$ as functions and operators on the larger set $\mathcal{Y}$ (and vice-versa). We define the restriction of $f$ to $\mathcal{S}$ as
$f|_{\mathcal{S}}(x)\coloneqq f(x)\text{ for all $x\in\mathcal{S}$}.$
For any $g\in\mathbb{R}^{\mathcal{S}}$ we define $g^{\uparrow\mathcal{Y}}\in\mathbb{R}^{\mathcal{Y}}$, the extension of $g$ to $\mathcal{Y}$, as
$\big[g^{\uparrow\mathcal{Y}}\big](x)\coloneqq\begin{cases}g(x)&\quad\text{ if $x\in\mathcal{S}$},\\ 0&\quad\text{ if $x\in\mathcal{Y}\setminus\mathcal{S}$}.\end{cases}$
This extension is the right inverse of restriction, namely
$\big(g^{\uparrow\mathcal{Y}}\big)\big|_{\mathcal{S}}=g,$
for all $g\in\mathbb{R}^{\mathcal{S}}$.
Let $M:\mathbb{R}^{\mathcal{Y}}\to\mathbb{R}^{\mathcal{Y}}$ . Define the restriction of $M$ to $\mathcal{S}$ as
$M\big|_{\mathcal{S}}\,g\coloneqq\left(M\big(g^{\uparrow\mathcal{Y}}\big)\right)\Big|_{\mathcal{S}}\ \text{ for all $g\in\mathbb{R}^{\mathcal{S}}$}.$
If $M$ is a linear mapping, i.e. a $|\mathcal{Y}|\times|\mathcal{Y}|$ matrix, then $M|_{\mathcal{S}}$ coincides with the $|\mathcal{S}|\times|\mathcal{S}|$ submatrix indexed by $\mathcal{S}$.
2.4 Closed-Form Solution for Hitting Probabilities
In this section, we derive a closed-form expression for hitting probabilities for homogeneous Markov chains. Although most of the material in this section appears in standard references (e.g. [norrisbook1997markov, kemeny1976finite, DoyleSnell1984, GrinsteadSnelll1997]), we present it here in full detail, since the terminology and results will be essential for the subsequent development in the imprecise setting.
By Proposition 2.5, a homogeneous Markov chain $\mathbb{P}_{T}$ with transition matrix $T$, hitting probabilities are trivial for states in $A\cup\mathcal{C}_{T}$. The system of equations in (4), as we restrict to $A^{c}\setminus\mathcal{C}_{T}$, can be reduced to
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=(Tp^{T})|_{A^{c}\setminus\mathcal{C}_{T}},$ (5)
where $p^{T}$ denotes the vector of hitting probabilities of $\mathbb{P}_{T}$. The following lemma, whose proof can be found in the appendix, allows us to rewrite the right-hand side of (5) in a more convenient way.
Lemma 2.6.
Let $\mathcal{S}\subset A^{c}\subset\mathcal{X}$, let $T$ be a transition matrix, and let $f\in\mathbb{R}^{\mathcal{X}}$ be any function such that $f(x)=1$ for all $x\in A$ and $f(x)=0$ for all $x\in\mathcal{S}$. Then
$(Tf)|_{A^{c}\setminus\mathcal{S}}=\bigl(T|_{A^{c}\setminus\mathcal{S}}\,(f|_{A^{c}\setminus\mathcal{S}})\bigr)+(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{S}}.$ (6)
This lemma allows us to rewrite the system of equations in (5) as
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(T|_{A^{c}\setminus\mathcal{C}_{T}}\,p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)+(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}}$
$\ \Rightarrow\ (I-T|_{A^{c}\setminus\mathcal{C}_{T}})p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}}.$
Hence, if we can show that the matrix $(I-T|_{A^{c}\setminus\mathcal{C}_{T}})$ is invertible, we can uniquely characterise $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ and have a closed-form expression to compute it. The structure of the argument is similar to the reasoning used in Krak [krak2021comphit, Lemma 10].
We proceed with the following lemma, the proof of which can be found in the appendix.
Lemma 2.7.
Let $T$ be a transition matrix. Fix any $x\in A^{c}\setminus\mathcal{C}_{T}$ and $n\in\mathbb{N}$ with $n>1$. Then if $T^{n}\mathbb{1}_{A}](x)>0$ and $T^{m}\mathbb{1}_{A}](x)=0$ for all $m\in\mathbb{N}$ with $m<n$, there exists some $y\in A^{c}\setminus\mathcal{C}_{T}$, $y\neq x$ such that $T(x,y)>0$ and $T^{n-1}\mathbb{1}_{A}](y)>0$.
Let $\mathbf{0}$ and $\mathbf{1}$ be the vectors made of all zeros and ones [^3] The size of these vectors will be clear for the context., respectively, and let $\mathcal{S}\subseteq\mathcal{X}$. Then, the quantity $(T|_{\mathcal{S}})^{k}\mathbf{1}](x)$, for $x\in\mathcal{S}$ and $k\in\mathbb{N}_{0}$, measures the probability that a homogeneous Markov chain with transition matrix $T$ starting in state $x$ stays in $\mathcal{S}$ after $k$ steps. Hence, it is not very surprising that this quantity is bounded below by zero and above by one.
Lemma 2.8.
Let $T$ be a transition matrix and consider an arbitrary subset of states $\mathcal{S}\subseteq\mathcal{X}$. Let $f,g\in\mathbb{R}^{\mathcal{S}}$ be such $f\leq g$. Then $(T|_{\mathcal{S}})^{k}f\leq(T|_{\mathcal{S}})^{k}g$ for all $k\in\mathbb{N}_{0}$. Moreover, we have that
$\mathbf{0}\leq(T|_{\mathcal{S}})^{k}\mathbf{1}\leq\mathbf{1}$ (7)
for all $k\in\mathbb{N}_{0}$.
The proof of this lemma can be found in the appendix.
Next, we show that for the choice $\mathcal{S}=A^{c}\setminus\mathcal{C}_{T}$, the upper bound in (7) becomes strict for all sufficiently large $k\in\mathbb{N}$. Indeed, by definition, every state in $A^{c}\setminus\mathcal{C}_{T}$ can reach $A$ in finitely many steps. Hence, if $k$ exceeds the minimal number of steps $n$ required for $x$ to reach a state outside $A^{c}\setminus\mathcal{C}_{T}$, the probability of remaining in $A^{c}\setminus\mathcal{C}_{T}$ after $k\geq n$ steps is strictly lower than $1$. The result below formalises this; its proof can be found in the appendix.
Lemma 2.9.
Let $T$ be a transition matrix, and let $A$ be the target set. Then there exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
$(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)<1\quad\text{for all }x\in A^{c}\setminus\mathcal{C}_{T}.$
Since the intuition for Lemma 2.9 was that it holds because every state in $A^{c}\setminus\mathcal{C}_{T}$ will reach a state outside of $A^{c}\setminus\mathcal{C}_{T}$, the result actually holds for every subset of $A^{c}\setminus\mathcal{C}_{T}$ as well. This slightly more general result will prove useful in the imprecise setting later on.
Corollary 2.10.
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists $n\in\mathbb{N}$ such that for all $k\in\mathbb{N}$ with $k\geq n$, it holds that
$(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)<1\quad\text{for all }x\in A^{c}\setminus\mathcal{S}.$
The proof of this corollary can be found in the appendix.
Using this result, we can conclude that powers of the transition matrix $T$ restricted to $A^{c}\setminus\mathcal{C}_{T}$ are smaller in norm than 1.
Corollary 2.11.
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then there exists some $n\in\mathbb{N}$ such that
$||(T|_{A^{c}\setminus\mathcal{S}})^{n}||<1,$
where $||\cdot||$ is the operator norm induced by the sup norm:
[…]
The proof of this result can be found in the appendix.
Corollary 2.11 allows us to finally obtain the results we are after regarding the invertibility of $(I-T|_{A^{c}\setminus\mathcal{C}_{T}})$.
Lemma 2.12.
Let $T$ be a transition matrix. Consider an arbitrary subset of states $\mathcal{S}\subset\mathcal{X}$ such that $\mathcal{C}_{T}\subseteq\mathcal{S}\subset A^{c}$. Then
$\lim_{k\to+\infty}(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}=\mathbf{0}$
and the inverse $(I-T|_{A^{c}\setminus\mathcal{S}})^{-1}$ exists and satisfies
$(I-T|_{A^{c}\setminus\mathcal{S}})^{-1}=\sum_{k=0}^{\infty}(T|_{A^{c}\setminus\mathcal{S}})^{k}.$
Proof.
By Corollary 2.11, there exists some $n\in\mathbb{N}$ such that $||(T|_{A^{c}\setminus\mathcal{S}})^{n}||<1$. This means that the restriction of $T$ to $A^{c}\setminus\mathcal{S}$ eventually becomes a contraction and the probability of remaining there decays to zero. Hence, the Neumann series $\sum_{k=0}^{\infty}(T|_{A^{c}\setminus\mathcal{S}})^{k}$ converges and the result follows (Kemeny and Snell [kemeny1976finite] give more details). ∎
The next proposition gives a closed-form expression for the vector of hitting probabilities of a given transition matrix $T$.
Proposition 2.13.
Let $\mathbb{P}_{T}$ be a homogeneous Markov chain defined on the state space $\mathcal{X}$ with transition matrix $T$. Then $p^{T}(x)=1$ for $x\in A$ and $p^{T}(x)=0$ if and only if $x\in\mathcal{C}_{T}$. Furthermore, the vector $p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}$ is given by
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(I-T|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)^{-1}(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}}.$ (8)
Proof.
The first two claims directly follow from Proposition 2.5. Restricting the system in Equation (4) to $A^{c}\setminus\mathcal{C}_{T}$ yields
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=(Tp^{T})|_{A^{c}\setminus\mathcal{C}_{T}}.$
By Lemma 2.6, we find that
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(T|_{A^{c}\setminus\mathcal{C}_{T}}\,p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)+(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}}.$
Hence,
$p^{T}|_{A^{c}\setminus\mathcal{C}_{T}}=\bigl(I-T|_{A^{c}\setminus\mathcal{C}_{T}}\bigr)^{-1}(T\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{C}_{T}},$
where we applied Lemma 2.12 to guarantee the existence of the inverse. ∎
The last lemma we present in this section is required for results in the imprecise setting later on. Given a transition matrix $T$, consider the induced directed weighted graph containing as vertices the states of $\mathcal{X}$ and directed edges with weight $T(x,y)$ for all $x,y\in\mathcal{X}$. Then, the following lemma establishes that, from any state $x\in\mathcal{X}$ which can reach $A$, there must be a (directed) simple path from $x$ to $A$ using only nonzero weight edges for which the hitting probabilities of the states along the way are nondecreasing as we approach $A$. Essentially, this says that from every such state $x$, there must be a way to “get closer” (that is, increase the hitting probability) to $A$.
Lemma 2.14.
Consider a homogeneous Markov chain $\mathbb{P}_{T}$ with transition matrix $T$ and let $p^{T}$ be its vector of hitting probabilities. For all $x\in A^{c}\setminus\mathcal{C}_{T}$, there exists a simple path $x=x_{0},x_{1},\dots,x_{n}=z\in A$ with $T(x_{k},x_{k+1})>0$ and
$p^{T}(x_{k})\leq p^{T}(x_{k+1})\quad\text{for all }k\in\{0,\dots,n-1\}.$ (9)
The proof of this lemma can be found in the appendix.
2.5 Imprecise Markov Chains
Instead of a single stochastic process $\mathbb{P}$, we now consider a set of stochastic processes $\mathcal{P}$. If these processes are Markov chains, it is said to be an imprecise Markov chain (IMC). An interpretation may be that we do not know precisely which $\mathbb{P}\in\mathcal{P}$ is the true process so that the set $\mathcal{P}$ serves as an “error margin”, and, we hope, contains the true process $\mathbb{P}$ being modelled as well. Inferences with respect to the imprecise Markov chain $\mathcal{P}$ are defined using lower and upper expectations, which are defined as
$\underline{\mathbb{E}}_{\mathcal{P}}[\cdot\mid\cdot]\coloneqq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\cdot\mid\cdot]\quad\text{and}\quad\overline{\mathbb{E}}_{\mathcal{P}}[\cdot\mid\cdot]\coloneqq\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\cdot\mid\cdot].$
These inferences are conservative, tight lower and upper bounds on inferences with respect to all stochastic processes in the set $\mathcal{P}$. Lower and upper probabilities, $\underline{\mathbb{P}}_{\mathcal{P}}$ and $\overline{\mathbb{P}}_{\mathcal{P}}$, are a special case of lower and upper expectations:
$\underline{\mathbb{P}}_{\mathcal{P}}(E\mid\cdot)\coloneqq\underline{\mathbb{E}}_{\mathcal{P}}[\mathbb{1}_{E}\mid\cdot]\quad\text{and}\quad\overline{\mathbb{P}}_{\mathcal{P}}(E\mid\cdot)\coloneqq\overline{\mathbb{E}}_{\mathcal{P}}[\mathbb{1}_{E}\mid\cdot],$
for all events $E\in\mathcal{F}$.
The set $\mathcal{P}$ representing the IMC will be parametrised through a nonempty set of transition matrices $\mathcal{T}$. Here we will focus only on (time-)homogeneous Markov chains $\mathbb{P}_{T}\in\mathcal{P}$ that have their transition matrix $T\in\mathcal{T}$. This choice is justified by the fact that Krak et al. [krak2019hitting] showed that lower (and upper) hitting probabilities remain unchanged even when the set $\mathbb{P}$ is enlarged to include (time-)inhomogeneous Markov processes compatible with $\mathcal{T}$.
Given a target set $A\subset\mathcal{X}$, we define the lower and upper hitting probabilities of the imprecise Markov chain parametrised by $\mathcal{T}$ as
$\underline{p}_{\mathcal{T}}\coloneqq\inf_{T\in\mathcal{T}}p^{T}\quad\text{and}\quad\overline{p}^{\mathcal{T}}\coloneqq\sup_{T\in\mathcal{T}}p^{T}.$
Hence, we have that
$\underline{p}^{\mathcal{T}}(x)=\underline{\mathbb{P}}_{\mathcal{P}}\Big(\exists n\in\mathbb{N}:X_{n}\in A\mid X_{0}=x\Big),$
and
$\overline{p}^{\mathcal{T}}(x)=\overline{\mathbb{P}}_{\mathcal{P}}\Big(\exists n\in\mathbb{N}:X_{n}\in A\mid X_{0}=x\Big).$
For a (precise) homogeneous Markov chain, the corresponding transition matrix $T$ provides valuable computational tools. We introduce analogous objects for imprecise Markov chains: the lower and upper transition operators.
Definition 2.15 (Lower and Upper Transition Operators [DeCooman2009]).
An operator $\underline{T}$ (resp. $\overline{T}$) is a lower (resp. upper) transition operator if, for all $f,g\in\mathbb{R}^{\mathcal{X}}$, all $\lambda>0$, and all $x\in\mathcal{X}$, it satisfies
  1. 1.
    $\min\limits_{y\in\mathcal{X}}f(y)\leq[\underline{T}f](x)$ and $\overline{T}f](x)\leq\max\limits_{y\in\mathcal{X}}f(y),$
  2. 2.
    $\underline{T}f+\underline{T}g\leq\underline{T}(f+g)$ and $\overline{T}f+\overline{T}g\geq\overline{T}(f+g),$
  3. 3.
    $\underline{T}(\lambda f)=\lambda\underline{T}f$ and $\overline{T}(\lambda f)=\lambda\overline{T}f.$
We note that lower and upper transition operators are conjugate, meaning that any lower transition operator $\underline{T}$ induces a corresponding upper transition operator
$\overline{T}(\cdot)=-\underline{T}(-\cdot),$ (10)
and vice-versa. Furthermore, any transition matrix is a lower transition operator with itself as the conjugate upper transition operator. Therefore, all properties derived for lower and upper transition operators also hold for transition matrices.
Beyond the properties in Definition 2.15, transition operators satisfy some additional useful properties.
Lemma 2.16 ([krak2019hitting]).
Given any lower transition operator $\underline{T}$ and its conjugate upper transition operator $\overline{T}$, for all $f,g\in\mathbb{R}^{\mathcal{X}}$ and all $n\in\mathbb{N}$ it holds that $\underline{T}^{n}$ is a lower transition operator, $\overline{T}^{n}$ its conjugate, and
  1. 1.
    $f\leq g\Rightarrow{\underline{T}^{n}}f\leq{\underline{T}^{n}}g$ and ${\overline{T}^{n}}f\leq{\overline{T}^{n}}g,$
  2. 2.
    ${\underline{T}^{n}}f+{\underline{T}^{n}}g\leq{\underline{T}^{n}}(f+g)$ and $\overline{T}^{n}(f+g)\leq\overline{T}^{n}f+\overline{T}^{n}g,$
  3. 3.
[…]
Fix some nonempty set of transition matrices $\mathcal{T}$. We will define the lower and upper transition operators corresponding to $\mathcal{T}$ as the lower and upper envelopes of $\mathcal{T}$, respectively.
Definition 2.17 ([krak2019hitting]).
Given a nonempty set of transition matrices $\mathcal{T}$, the corresponding lower and upper transition operators $\underline{T}$ and $\overline{T}$ are defined as
$\underline{T}f](x)\coloneqq\inf_{T\in\mathcal{T}}[Tf](x)\quad\text{and}\quad[\overline{T}f](x)\coloneqq\sup_{T\in\mathcal{T}}[Tf](x),$
for all $f\in\mathbb{R}^{\mathcal{X}}$ and all $x\in\mathcal{X}$.
Throughout the paper, we assume that $\mathcal{T}$ is nonempty, compact, convex, and with separately specified rows (SSR) [HermansSkulj2014, krak2019hitting]. Equivalently, $\mathcal{T}$ can be represented as $N$ compact and convex sets of probability distributions $\{\mathcal{T}_{x}\}_{x\in\mathcal{X}}$, one for each state $x\in\mathcal{X}$, so $\mathcal{T}$ is the Cartesian product of these sets. Each of these sets is often called credal set.
Proposition 2.18 ([krak2019hitting]).
Let $\mathcal{T}$ be a nonempty, compact, convex set of transition matrices with SSR on the space $\mathcal{X}$. Then, for all $f\in\mathbb{R}^{\mathcal{X}}$, there exists an extreme point $T\in\mathcal{T}$ such that $Tf=\underline{T}f$, and there exists an extreme point $T^{\prime}\in\mathcal{T}$ such that $T^{\prime}f=\overline{T}f$.
3 Characterization of Lower and Upper Hitting Probabilities
In this section, we characterise lower and upper hitting probabilities as the unique solutions of certain systems of equations. We first formalise the notion of reachability in the imprecise framework and use it to identify the states that can reach the target set with positive lower (resp. upper) probability.
3.1 Lower and Upper Reachability
Before characterising the lower and upper hitting probabilities, we identify the subset of states for which these probabilities are trivial, i.e. equal to zero or one. Recall that in the precise setting, for a homogeneous Markov chain with transition matrix $T$, we focused on the set of states $A^{c}\setminus\mathcal{C}_{T}$. This was done because Proposition 2.5 directly yields the hitting probabilities for those states, with the hitting probability equal to one for states in $A$ and equal to zero for states in $\mathcal{C}_{T}$. Clearly, the result for the states in $A$ carries over to the imprecise setting, as $p^{T}(x)=1$ for all $T\in\mathcal{T}$ and all $x\in A$. Hence also $\underline{p}^{\mathcal{T}}(x)=\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$. To characterise the states with zero lower and upper hitting probability, we need suitable definitions of reachability for the imprecise setting.
Definition 3.1 (Upper reachability [DeCooman2009, hermans2012characterisation, debock2017imprecise]).
Given a set of transition matrices $\mathcal{T}$ we say that a set of states $\mathcal{D}\subseteq\mathcal{X}$ is upper reachable from $x\in\mathcal{X}$ if there exists $n\in\mathbb{N}_{0}$ such that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$, i.e. there exists $n\in\mathbb{N}_{0}$ and $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoonup\mathcal{D}$ if $\mathcal{D}$ is upper reachable from $x$.
The fact that $x\rightharpoonup\mathcal{D}$ if and only if there exists $T\in\mathcal{T}$ and $n\in\mathbb{N}_{0}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$, follows from the fact that $\overline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ implies the existence of a sequence of matrices $T_{1},\dots,T_{n}\in\mathcal{T}$ such that
$T_{1}\cdots T_{n}\mathbb{1}_{\mathcal{D}}](x)>0.$
From this sequence, and using the fact that the set $\mathcal{T}$ is convex and has SSR, we construct the desired matrix $T$.
Definition 3.2 (Lower reachability $\operatorname{\mathbf{(LR0)}}$).
Given a set of transition matrices $\mathcal{T}$ we say that a set of states $\mathcal{D}\subseteq\mathcal{X}$ is lower reachable from $x\in\mathcal{X}$ if for all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}_{0}$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. We denote by $x\rightharpoondown\mathcal{D}$ if $\mathcal{D}$ is lower reachable from $x$.
The definition of upper reachability is widely accepted meanwhile the same does not hold for the lower reachability one. In fact, in the literature [debock2025convergent, hermans2012characterisation, sangalli2025meeting, debock2017imprecise, erreygers2023], at least three other definitions are used.
(LR1):
$x$ lower reaches $\mathcal{D}$ if $x\in\mathcal{D}_{k^{*}}$ where $(\mathcal{D}_{k})_{k\in\mathbb{N}_{0}}$ is a nondecreasing sequence given by $\mathcal{D}_{0}=\mathcal{D}$ and
$\mathcal{D}_{k+1}=\mathcal{D}_{k}\cup\{z\in\mathcal{X}\setminus\mathcal{D}_{k}\mid[\underline{T}\mathbb{1}_{\mathcal{D}_{k}}](z)>0\},$
and where $k^{*}$ is the first index for which $\mathcal{D}_{k^{*}}=\mathcal{D}_{k^{*}+1}$. [debock2017imprecise, erreygers2023]
(LR2):
$x$ lower reaches $\mathcal{D}$ if there exists $n\in\mathbb{N}_{0}$ such that for all $T\in\mathcal{T}$ it holds that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. [sangalli2025meeting]
(LR3):
$x$ lower reaches $\mathcal{D}$ if $\underline{T}^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for some $n\in\mathbb{N}_{0}$. [debock2025convergent, sangalli2025meeting]
Let us note that one reason for this abundance of possible definitions of lower reachability is that the quantifiers do not commute, i.e. in general
$\exists n\in\mathbb{N}_{0}:\forall T\in\mathcal{T},\text{(property)}$” and $\forall T\in\mathcal{T},\exists n\in\mathbb{N}_{0}:\text{(property)}$
are not equivalent. Conversely, for upper reachability the quantifiers $\exists n\in\mathbb{N}_{0}$” and $\exists T\in\mathcal{T}$” commute: each formulation just requires the existence of at least one suitable pair $(n,T)\in\mathbb{N}_{0}\times\mathcal{T}$.
Theorem 3.3.
Definition $\operatorname{(LR0)}$ and definition $\operatorname{(LR1)}$ are equivalent.
Proof.
$\operatorname{(LR0)}\Rightarrow\operatorname{(LR1)}:$ Suppose ex-absurdo that $x\notin\mathcal{D}_{k^{*}}$. This implies that there exists $\tilde{T}\in\mathcal{T}$ such that $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$.
Let $x=x_{0},\dots,x_{n}=y\in\mathcal{D}$ be a simple path from Proposition 2.2 that connects $x$ to $y$ using $\tilde{T}$. This path exists because $x\rightharpoondown\mathcal{D}$ under $\operatorname{(LR0)}$. Since $\tilde{T}(x,z)=0$ for all $z\in\mathcal{D}_{k^{*}}$ and $\tilde{T}(x,x_{1})>0$, it follows that $x_{1}\notin\mathcal{D}_{k^{*}}$. Note that the condition $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$ involves only the $x$th row of the matrix $\tilde{T}$. Therefore, since $\mathcal{T}$ has SSR and $x_{1}\notin\mathcal{D}_{k^{*}}$, there exists a modification of the $x_{1}$th row of $\tilde{T}$ such that $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{1})=0$. We then proceed along the path: if $\tilde{T}(x_{1},x_{2})>0$ we find that $x_{2}\notin\mathcal{D}_{k^{*}}$, and so on. Thus, we either find $x_{j}\notin\mathcal{D}_{k^{*}}$, for all $0\leq j\leq n$, which is in contradiction with $x_{n}\in\mathcal{D}\subseteq\mathcal{D}_{k^{*}}$, or we find an index $j^{*}$ such that by imposing $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{j^{*}})=0$ we get $\tilde{T}(x_{j^{*}},x_{j^{*}+1})=0$. This means that $x_{j}\notin\mathcal{D}_{k^{*}}$, for all $0\leq j\leq j^{*}$, while $x_{j^{*}+1}$ is in $\mathcal{D}_{k^{*}}$ and, for that reason, $\tilde{T}(x_{j^{*}},x_{j^{*}+1})$ has been set to zero.
If $x_{j^{*}}\not\to^{\tilde{T}}\mathcal{D}$ with this modified $\tilde{T}$, we restart from $x$. By our hypothesis, definition $\operatorname{(LR0)}$, and Lemma 2.14, there exists $m\in\mathbb{N}$, a matrix $\hat{T}\in\mathcal{T}$, with $\hat{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](x_{j})=0$ for all $1\leq j\leq j^{*}$, and a path $x=y_{0},\dots,y_{m}\in\mathcal{D}$ such that $0<p^{\hat{T}}(x)\leq\dots\leq p^{\hat{T}}(y_{m})=1$. This path cannot contain $x_{j^{*}}$ since its hitting probability is zero. As before, we impose $\hat{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](y_{i})=0$ along the path until we either reach $\mathcal{D}$ and get a contradiction or we encounter the first state $y_{i^{*}}$ for which $\hat{T}(y_{i^{*}},y_{i^{*}+1})=0$. If $y_{i^{*}}$ cannot reach $\mathcal{D}$ we eliminate $y_{i^{*}}$ and restart from $x$: the finiteness of $\mathcal{X}$ guarantees that we may repeat this argument only a finite number of times.
If $y_{i^{*}}\to^{\hat{T}}\mathcal{D}$, let $y_{i^{*}}=w_{0},\dots,w_{l}\in\mathcal{D}$ be a simple path that connects $y_{i^{*}}$ to $\mathcal{D}$ using this modified version of $\hat{T}$. Again, we follow this path until we get a contradiction or we encounter the first index $h^{*}$ for which by imposing $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](w_{h^{*}})=0$ we get $\tilde{T}(w_{h^{*}},w_{h^{*}+1})=0$. We argue that, since $w_{h^{*}+1}\in\mathcal{D}_{k^{*}}$, $w_{h^{*}}$ is not in $\{x_{0},\dots,x_{j^{*}}\}$ or $\{y_{0},\dots,y_{i^{*}}\}$. Indeed, for these states, we have already set $\tilde{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](\cdot)=0$, so the path $y_{i^{*}}=w_{0},\dots,w_{l}$ that included one of these states could not have had positive transition probabilities. Therefore, each time we build a path we discover a new state not in $\mathcal{D}_{k^{*}}$: using the finiteness of $\mathcal{X}$ we get a contradiction.
$\operatorname{(LR1)}\Rightarrow\operatorname{(LR0)}:$ If $x\in\mathcal{D}_{1}\setminus\mathcal{D}$ then, by the definition of the sequence $(\mathcal{D}_{k})$, $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$ for all $T\in\mathcal{T}$, with $n=1$.
We will now proceed by strong induction on the following proposition:
$(P_{k}):$
If $x\in\mathcal{D}_{k}\setminus\mathcal{D}_{k-1}$, then for all $T\in\mathcal{T}$ there exists $n\leq k$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$.
Suppose that for all $w\in\mathcal{D}_{j}\setminus\mathcal{D}_{j-1}$ and all $1\leq j<k$ we have that $(P_{j})$ is true, i.e. there exists a path of length $m\leq j$ connecting $w$ to $\mathcal{D}$. This implies that for all $w\in\mathcal{D}_{k-1}$ and for all $T\in\mathcal{T}$ there exists a path of length $m\leq k-1$ connecting $w$ to $\mathcal{D}$, i.e.
${T}^{m}\mathbb{1}_{\mathcal{D}}](w)=\sum_{\begin{subarray}{c}i=2,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(w,x_{2})\cdots{T}(x_{m},x_{m+1})>0.$
For all $x\in\mathcal{D}_{k}\setminus\mathcal{D}_{k-1}$ and all $T\in\mathcal{T}$ it holds that
$T\mathbb{1}_{\mathcal{D}_{k-1}}](x)=\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)>0.$
It follows that for all $T\in\mathcal{T}$ there exists $m\leq k-1$ such that
$[{T}^{m+1}\mathbb{1}_{\mathcal{D}}](x)$ $=\sum_{\begin{subarray}{c}i=1,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(x,x_{1})\cdots{T}(x_{m},x_{m+1})$
$\geq\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)\sum_{\begin{subarray}{c}i=2,\dots,m\\ x_{i}\in\mathcal{X}\\ x_{m+1}\in\mathcal{D}\end{subarray}}{T}(w,x_{2})\cdots{T}(x_{m},x_{m+1})$
$=\sum_{w\in\mathcal{D}_{k-1}}{T}(x,w)[{T}^{m}\mathbb{1}_{D}](w)>0,$
which is what we wanted to prove. ∎
In the rest of the paper we will use (LR0) (Definition 3.2) (or the equivalent (LR1)) to define lower reachability. (LR0) is weaker than (LR2) and (LR3), namely
$\text{(LR3)}\Rightarrow\text{(LR2)}\Rightarrow\text{(LR1)}\Leftrightarrow\text{(LR0)}.$ (11)
The following examples show why these implications cannot be reversed.
Example 3.4.
This example shows that $\operatorname{(LR0)}\not\Rightarrow\operatorname{(LR2)}$. Let $\mathcal{X}=\{1,2,3,4\}$ be the state space, and let $\{3\}$ be the target set. Let the set of transition matrices on $\mathcal{X}$ be
$\mathcal{T}=\left\{T=\begin{bmatrix}0&p&1-p&0\\ 0&0&1&0\\ 0&0&0&1\\ 0&0&0&1\end{bmatrix}:p\in[0,1]\right\}.$
The situation is represented by the following graph.
1234pp1p1-p
For all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}$ (either 1 or 2) such that $T^{n}(1,3)>0$. Hence, $3$ is lower reachable from $1$, under $\operatorname{(LR0)}$. However, it is not true that there exists $n\in\mathbb{N}$ such that $T^{n}(1,3)>0$ for all $T\in\mathcal{T}$. In fact, the matrix $T_{1}\in\mathcal{T}$, defined by fixing $T_{1}(1,2)=p=1$, has $T_{1}^{2}(1,3)=1>0$ and $T_{1}^{1}(1,3)=0$. On the other hand, the matrix $T_{0}\in\mathcal{T}$, defined by fixing $T_{0}(1,2)=p=0$, has $T_{0}^{1}(1,3)=1$ and $T_{0}^{2}(1,3)=0$. So $3$ is not lower reachable from $1$, under $\operatorname{(LR2)}$, because for all $n\in\mathbb{N}_{0}$ there exists $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\{3\}}](1)=0$.
Example 3.5.
This example shows that $\operatorname{(LR2)}\not\Rightarrow\operatorname{(LR3)}$. Let $\mathcal{X}=\{1,2,3\}$ be the state space, and let $\{3\}$ be the target set. Let the set of transition matrices on $\mathcal{X}$ be
$\mathcal{T}=\left\{T=\begin{bmatrix}0&1-q&q\\ 0&0&1\\ 1&0&0\\ \end{bmatrix}:q\in[0,1]\right\}.$
The situation is represented by the following graph:
1231q1-qqq
If $n=6$, all $T\in\mathcal{T}$ are such that $T^{6}(1,3)>0$, i.e. in six steps, whichever $T$ is selected, the process has a positive probability of being in state 3, if it has started in state 1. Therefore, 3 is lower reachable from 1, under $\operatorname{(LR2)}$. However, for all $n\in\mathbb{N}$ we have
$\underline{T}^{n}\mathbb{1}_{\{3\}}](1)=0,$
implying that 3 is not lower reachable from 1, under $\operatorname{(LR3)}$. This comes from the fact that, since we are applying the lower transition operator $\underline{T}$ $n$ times, there is no guarantee that the same transition matrix achieves the minimum. Given $n\in\mathbb{N}$, we can find a sequence of $n$ transition matrices that minimises the probability of hitting state 3 in exactly $n$ steps, and we find that this probability is zero. Given $n\in\mathbb{N}$, the sequence of transition matrices $(T_{k})_{k\in\mathbb{N}}$ defined as
$T_{k}\coloneqq\begin{cases}S_{0},&\text{ if $n$ is odd and $k=1$},\\ S_{1},&\text{otherwise},\end{cases}$
where $S_{0,1}\in\mathcal{T}$ are the two matrix defined by setting either $q=0$ or $q=1$, ensures to avoid state 3 at time $n$. Thus, 3 is not lower reachable from 1, under $\operatorname{(LR3)}$.
There are particular cases in which (LR0), (LR1), (LR2) and (LR3) are all equivalent. We say that the set $\mathcal{D}$ is closed when for all $x\in\mathcal{D},y\notin\mathcal{D}$ and for all $T\in\mathcal{T}$, we have $T(x,y)=0$.
Proposition 3.6.
Under the assumption that $\mathcal{D}$ is closed, definitions $\operatorname{(LR0)}$, $\operatorname{(LR1)}$, $\operatorname{(LR2)}$ and $\operatorname{(LR3)}$ are equivalent.
Proof.
The fact that (LR0) $\Rightarrow$ (LR2) comes from the fact that (LR0) guarantees for all $T\in\mathcal{T}$ the existence of a time $n=n(T)$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. Then $n^{*}\coloneqq\max_{T\in\mathcal{T}}n(T)$ is the time for which (LR2) is satisfied.
To prove that $x$ lower reaches $y$ under (LR3) if it does under (LR1), let us recall that we need to prove that there exists $n\in\mathbb{N}$ such that
$\sum_{x_{1},\dots,x_{n-1}\in\mathcal{X}}T_{1}(x,x_{1})\cdots T_{n}(x_{n-1},y)>0,$
for all $T_{1},\dots,T_{n}\in\mathcal{T}$. Let us pick $n=k^{*}$ recalling that, by hypothesis, we have $x\in\mathcal{D}_{k^{*}}$. For all $T_{1}\in\mathcal{T}$, there exists $x_{1}\in\mathcal{D}_{k^{*}-1}$ such that $T_{1}(x,x_{1})>0$. Iterating this argument will get to $\mathcal{D}$ in at most $k^{*}$ steps. If, for some sequence of matrices, there is a positive probability of hitting $\mathcal{D}$ at time $k^{\prime}<k^{*}$, then, due to the closedness of $\mathcal{D}$, the same would hold at time $k^{*}$. ∎
3.2 Lower Hitting Probabilities
Having discussed the definition of reachability extensively, we can now use this definition to characterise lower hitting probabilities, i.e. $\underline{p}^{\mathcal{T}}=\inf_{T\in\mathcal{T}}p^{T}$.
Definition 3.7.
We define the set of states $\mathcal{A}_{\mathcal{T}}\subset\mathcal{X}$ that cannot lower reach the target as
$\mathcal{A}_{\mathcal{T}}\coloneqq\{x\in A^{c}\mid x\not\rightharpoondown A\}.$ (12)
This set can be rewritten as
$\mathcal{A}_{\mathcal{T}}$ $=\{x\in A^{c}\mid\exists T\in\mathcal{T}:\forall n\in\mathbb{N},\ [T^{n}\mathbb{1}_{A}](x)=0\}$
$=\{x\in A^{c}\mid\exists T\in\mathcal{T}:x\in\mathcal{C}_{T}\}.$
We may interpret $\mathcal{A}_{\mathcal{T}}$ as the set of states whose lower probability of reaching $A$ is zero, since there exists a transition matrix that disconnects $\mathcal{A}_{\mathcal{T}}$ from $A$. Using the definition of lower reachability, we deduce a straightforward relationship between the sets $\mathcal{A}_{\mathcal{T}}$ and $\mathcal{C}_{T}$.
Lemma 3.8.
It holds that
$\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}.$
Proof.
The proof follows immediately from the definition of $\mathcal{A}_{\mathcal{T}}$ and lower reachability. ∎
We are finally able to characterise lower hitting probabilities as the unique solution to a system of equations.
Proposition 3.9.
The vector of lower hitting probabilities $\underline{p}^{\mathcal{T}}$ satisfies
1.
$\underline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$,
2.
$\underline{p}^{\mathcal{T}}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$,
3.
$0<\underline{p}^{\mathcal{T}}(x)\leq 1$ for all $x\in A^{c}\setminus\mathcal{A}_{\mathcal{T}}$,
and it is the only solution to
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\underline{T}\,\underline{p}^{\mathcal{T}}.$ (13)
satisfying conditions 1 and 2.
Proof.
This result is derived similarly to Krak’s Proposition 2 [krak2021comphit]. Let us start with the claims about the structure of $\underline{p}^{\mathcal{T}}$. We know from Krak et al. [krak2019hitting, Corollary 19] that Equation (13) has a minimal nonnegative solution $\underline{p}^{\mathcal{T}}\in[0,1]^{\mathcal{X}}$. By the definition of the system, this solution satisfies $\underline{p}^{\mathcal{T}}(x)=1$ for $x\in A$. The fact that $\underline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in\mathcal{A}_{\mathcal{T}}$ follows from the definition of lower hitting probability, Lemma 3.8, and Proposition 2.2.
Now we prove the uniqueness of $\underline{p}^{\mathcal{T}}$. We want to show that if $p_{1},p_{2}\in[0,1]^{\mathcal{X}}$ are solutions of the system in Equation (13) and have $p_{1}(x)=p_{2}(x)=1$ for all $x\in A$ and $p_{1}(x)=p_{2}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$, then it must be $p_{1}=p_{2}$. Note that it suffices to show that $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$.
By Propositions 2.18, there exist $T_{1},T_{2}\in\mathcal{T}$ such that
$T_{1}p_{1}=\underline{T}p_{1}\quad\text{and}\quad T_{2}p_{2}=\underline{T}p_{2}.$
Since $p_{1}$ and $p_{2}$ are solutions to the system, it holds that
$p_{1}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{1}p_{1},$ (14)
and
$p_{2}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{2}p_{2}.$ (15)
Since $p_{1}$ and $p_{2}$ satisfy the systems in Equations (14) and (15), respectively, and both satisfy (13), we find that
$(T_{1}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\quad\text{ and }\quad(T_{2}p_{2})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{2})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}.$
We then get
$(T_{1}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\leq(T_{2}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}.$
Equation (14) then yields
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ $=(T_{1}p_{1})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$\leq(T_{2}p_{1})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$=T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\,p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
where we applied Lemma 2.6 in the last step. Recursively expanding this expression for $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ yields, after $m\in\mathbb{N}$ expansions:
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{m}p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+\sum_{k=0}^{m-1}(T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{k}(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}.$
Observe that $\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}$ by Lemma 3.8 and $\mathcal{C}_{T_{2}}\subseteq\mathcal{A}_{\mathcal{T}}\subset A^{c}$. Taking the limits as $m\to+\infty$ and applying Lemma 2.12 yields
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(I-T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{-1}(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
where the last equality follows from Proposition 2.13. Therefore, without additional assumptions on $p_{1}$ and $p_{2}$, we have $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$. Similarly, by symmetry, $p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$, and thus $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$. ∎
3.3 Upper Hitting Probabilities
We now turn our attention to characterising upper hitting probabilities, i.e. $\overline{p}^{\mathcal{T}}=\sup_{T\in\mathcal{T}}p^{T}$.
Definition 3.10.
We define the set of states $\mathscr{W}_{\mathcal{T}}$} that cannot upper reach the target as
${\mathscr{W}_{\mathcal{T}}}\coloneqq\{x\in A^{c}\mid x\not\rightharpoonup A\}.$ (16)
This set can be rewritten as
${\mathscr{W}_{\mathcal{T}}}$ $=\{x\in A^{c}\mid\forall T\in\mathcal{T},\ \forall n\in\mathbb{N}:[T^{n}\mathbb{1}_{A}](x)=0\}$
$=\{x\in A^{c}\mid\forall T\in\mathcal{T}:x\in\mathcal{C}_{T}\}.$
We may interpret $\mathscr{W}_{\mathcal{T}}$ as the set of states whose upper probability of reaching $A$ is zero, since every transition matrix disconnects $\mathscr{W}_{\mathcal{T}}$ to $A$. Similarly to lower hitting probabilities, there is a straightforward connection between sets $\mathscr{W}_{\mathcal{T}}$ and $\mathcal{C}_{T}$
Lemma 3.11.
It holds that
${\mathscr{W}_{\mathcal{T}}}=\bigcap_{T\in\mathcal{T}}\mathcal{C}_{T}.$
Proof.
The proof follows immediately from the definition of $\mathscr{W}_{\mathcal{T}}$ and upper reachability. ∎
The following result characterises upper hitting probabilities as the unique solution to a system of equations.
Proposition 3.12.
The vector of upper hitting probabilities $\overline{p}^{\mathcal{T}}$ satisfies
I.
$\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$,
II.
$\overline{p}^{\mathcal{T}}(x)=0$ for all $x\in\mathscr{W}_{\mathcal{T}}$,
III.
$0<\overline{p}^{\mathcal{T}}(x)\leq 1$ for all $x\in A^{c}\setminus\mathscr{W}_{\mathcal{T}}$,
and it is the only solution to
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\overline{T}\,\overline{p}^{\mathcal{T}}$ (17)
satisfying conditions I and II.
Proof.
Let us start with the claims about the structure of $\overline{p}^{\mathcal{T}}$. We already know that Equation (17) has a minimal nonnegative solution $\overline{p}^{\mathcal{T}}\in[0,1]^{\mathcal{X}}$ [krak2019hitting, Corollary 19]. By the definition of the system, this solution satisfies $\overline{p}^{\mathcal{T}}(x)=1$ for $x\in A$. The fact that $\overline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in{\mathscr{W}_{\mathcal{T}}}$ follows from the definition of upper hitting probability and Lemma 3.11.
Now we prove the uniqueness of $\overline{p}^{\mathcal{T}}$. By the minimality of $\overline{p}^{\mathcal{T}}$ we have $\overline{p}^{\mathcal{T}}\leq p$ pointwise, for all $p\in[0,1]^{\mathcal{X}}$ solution to the system in Equation (17) with $p(x)=1$ for all $x\in A$ and $p(x)=0$ for all $x\in{\mathscr{W}_{\mathcal{T}}}$. Set
$d\coloneqq p-\overline{p}^{\mathcal{T}}\geq 0.$
Subtracting the two fixed-point equations yields
$d=\mathbb{1}_{A^{c}}\cdot\big(\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\big).$
Since $p=\overline{p}^{\mathcal{T}}+d$ and $d\geq 0$, subadditivity (Definition 2.15(2)) gives
$\overline{T}p=\overline{T}(\overline{p}^{\mathcal{T}}+d)\leq\overline{T}\overline{p}^{\mathcal{T}}+\overline{T}d.$
Hence $\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\leq\overline{T}d,$ and therefore
$d=\mathbb{1}_{A^{c}}\cdot\big(\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\big)\leq\mathbb{1}_{A^{c}}\cdot\overline{T}d.$
Applying this inequality $n$ times yields
$d\leq\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\mathbb{1}_{A^{c}}\cdot\overline{T}d\right)\leq\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\,\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\dots\left(\mathbb{1}_{A^{c}}\cdot\overline{T}d\right)\right)\right),$
which we write
$d\leq\left(\mathbb{1}_{A^{c}}\cdot\overline{T}\right)^{n}d.$
[…]
for all $n\in\mathbb{N}$. Therefore, we have
[…]
for all $n\in\mathbb{N}$.
Moreover, we observe that
$\Big(\left(\mathbb{1}_{A^{c}}\cdot\overline{T}\right)^{n}\mathbb{1}_{A^{c}}\Big)\Big|_{A^{c}}=\left(\overline{T}|_{A^{c}}\right)^{n}\mathbb{1}_{A^{c}}.$
For each $x\in A^{c}\setminus\mathscr{W}_{T}$ the quantity $((\overline{T}|_{A^{c}})^{n}\mathbb{1}_{A^{c}})(x)$ is the upper probability of avoiding $A$ for $n$ steps and therefore decays to 0 as $n$ approaches infinity. Taking the limit on both sides of (18) yields $d(x)\leq 0$ for every $x\in A^{c}\setminus\mathscr{W}_{\mathcal{T}}$, so $d=0$, which is what we wanted to prove. ∎
3.4 Further Discussions on the Definition of Reachability
An alternative and very natural definition of lower reachability can be formulated in terms of hitting probabilities:
(LRHP):
$x$ lower reaches $A$ if the lower probability of hitting $A$ is positive, i.e. $\underline{p}^{\mathcal{T}}(x)>0$.
Proposition 3.9 shows that this definition is equivalent to $\operatorname{(LR0)}$: indeed, $x\in\mathcal{A}_{\mathcal{T}}$, i.e. $x\not\rightharpoondown A$, if and only if $\underline{p}^{\mathcal{T}}(x)=0$. Taking the negation, we obtain the equivalence between $\operatorname{(LR0)}$ and $\operatorname{(LRHP)}$.
In the same way, the definition of upper reachability can be rewritten in terms of upper hitting probabilities.
4 Computing Hitting Probabilities for IMCs
Having established a characterisation of lower and upper hitting probabilities, this section presents practical algorithms for computing these quantities.
4.1 Computing Lower Hitting Probabilities
We show an algorithmic result for computing $\underline{p}^{\mathcal{T}}$. This result, which emulates Krak [krak2021comphit, Proposition 7], provides a practical algorithm for calculating $\underline{p}^{\mathcal{T}}$ iterating over extreme points of $\mathcal{T}$. The only condition is to start with a $T_{1}\in\mathcal{T}$ such that the trivial hitting probabilities are already correct (i.e. $p_{1}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$ and $p_{1}(x)=1$ for $x\in A$). Definition (LR1) provides a practical algorithm to identify the set $\mathcal{A}_{\mathcal{T}}$. In particular, if we start with $\mathcal{D}=A$ then $\mathcal{D}_{k^{*}}^{c}=\mathcal{A}_{\mathcal{T}}$. The matrix $T\in\mathcal{T}$ found in the last iteration of this algorithm is the desired starting matrix.
Proposition 4.1.
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the minimal nonnegative solution in $0,1]^{\mathcal{X}}$ to the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$
and let $T_{n+1}$ be an extreme point of $\mathcal{T}$ such that
$T_{n+1}p_{n}=\underline{T}p_{n}.$
Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the unique minimal solution of the system
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\underline{T}\,\underline{p}^{\mathcal{T}}.$ (19)
We delay the proof of this proposition until after the following lemma.
Lemma 4.2.
For the sequence of transition matrices $(T_{n})_{n\in\mathbb{N}}$ defined in Proposition 4.1 it holds that
$\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}},$ (20)
for all $n\in\mathbb{N}$.
Proof.
The proof uses induction on (20). For the base case $n=1$, we have $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$ by the choice of $T_{1}\in\mathcal{T}$. Assume now that $\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$ for some $n\in\mathbb{N}$. We will show that also $\mathcal{C}_{T_{n+1}}=\mathcal{A}_{\mathcal{T}}$. Let $x\in\mathcal{A}_{\mathcal{T}}$, then
$T_{n}p_{n}](x)=\sum_{y\in\mathcal{X}}T_{n}(x,y)p_{n}(y)=p_{n}(x)=0,$
since $\mathcal{A}_{\mathcal{T}}\subseteq A^{c}$ and $p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$. Note that the last equality holds by our induction hypothesis in combination with Proposition 2.5. By construction, we have
$T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n},$
since $T_{n}\in\mathcal{T}$. Therefore,
$0=[T_{n}p_{n}](x)\geq[T_{n+1}p_{n}](x)=\sum_{y\in\mathcal{X}}T_{n+1}(x,y)p_{n}(y).$
Since the right-hand side is nonnegative, we must have $T_{n+1}(x,y)=0$ for every $y\in\mathcal{X}$ with $p_{n}(y)>0$. By Proposition 2.5 this implies $T_{n+1}(x,y)=0$ for all $y\notin\mathcal{C}_{T_{n}}$. Applying the induction hypothesis, we obtain $T_{n+1}(x,y)=0$ for every $y\in\mathcal{A}_{\mathcal{T}}^{c}$, and hence $x\not\to^{T_{n+1}}y$ for all $y\in\mathcal{A}_{\mathcal{T}}^{c}$. Since $\mathcal{A}_{\mathcal{T}}\subseteq A^{c}$, it follows that $x\in\mathcal{C}_{T_{n+1}}$. As $x\in\mathcal{A}_{\mathcal{T}}$ was arbitrary, $\mathcal{A}_{\mathcal{T}}\subseteq\mathcal{C}_{T_{n+1}}$. Finally, using Lemma 3.8, which states $\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}$, we conclude that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n+1}}$, completing the induction. ∎
Proof of Proposition 4.1.
Let us first establish that the sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined. Since each $p_{n}$ is a vector of hitting probabilities, we have that $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$. Moreover, for any $n\in\mathbb{N}$, the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$
exists due to Proposition 2.5. Furthermore, for any $n\in\mathbb{N}$, the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.18.
Next, we will show that $(p_{n})_{n\in\mathbb{N}}$ is convergent and nonincreasing. By Proposition 2.5, we have that $p_{n}(x)=1$ for $x\in A$ for all $n\in\mathbb{N}$. Therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially nonincreasing and convergent on $A$. By Lemma 4.2, we have that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$, therefore by Proposition 2.5 it follows that $p_{n}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$ and all $n\in\mathbb{N}$, meaning that $(p_{n})_{n\in\mathbb{N}}$ is trivially nonincreasing and convergent on $\mathcal{A}_{\mathcal{T}}$.
Next, we will show that $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing on $A^{c}\setminus\mathcal{A}_{\mathcal{T}}$ as well. This is shown analogously to the proof of Krak’s Proposition 7 [krak2021comphit]. Consider any $n\in\mathbb{N}$. Recall that
$T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n}.$
This then yields that
$p_{n}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ $=(T_{n}p_{n})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$\geq(T_{n+1}p_{n})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$=T_{n+1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}p_{n}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+(T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
where we used Lemma 2.6 for the last step in combination with the fact that we have already shown that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yields that
$p_{n}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\geq(T_{n+1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{m}p_{n}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+\left[\sum_{k=0}^{m-1}(T_{n+1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{k}\right](T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}.$
Letting $m\to+\infty$ and applying Lemma 2.12 then yields that
$p_{n}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\geq(I-T_{n+1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{-1}(T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{n+1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
where the last step follows from Proposition 2.13. Thus the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing. Since $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists and lives in $0,1]^{{\mathcal{X}}}$.
Let us move on to showing that $p^{*}$ satisfies the system in Equation (19). Fix any $n\in\mathbb{N}$. Then
[…]
Continuing (21), we see that
[…]
where we used Lemma 2.16(3) for the second and third inequalities. Since all summands on the final right-hand side of (22) vanish as $n\to+\infty$, we find that $p^{*}$ solves the system in Equation (19).
Recall that $p_{n}\in[0,1]^{{\mathcal{X}}}$, $p_{n}(x)=1$ for $x\in A$ and $p_{n}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$ for all $n\in\mathbb{N}$. Hence also $p^{*}\in[0,1]^{{\mathcal{X}}}$, $p^{*}(x)=1$ for $x\in A$ and $p^{*}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$. Due to Proposition 3.9, we can then conclude that $p^{*}$ is the unique minimal solution to the system in Equation (19) and so $p^{*}=\underline{p}^{\mathcal{T}}$. ∎
4.2 Computing Upper Hitting Probabilities
We show an algorithmic result for computing the upper hitting probability $\overline{p}^{\mathcal{T}}$ similar to the one presented for $\underline{p}^{\mathcal{T}}$. Similarly to lower hitting probabilities, we need to initialise the algorithm with a matrix $T_{1}$ such that $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$. We can find such a matrix by taking any point in the interior of $\mathcal{T}$.
Proposition 4.3.
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}={\mathscr{W}_{\mathcal{T}}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$
and let $T_{n+1}$ be an extreme point of $\mathcal{T}$ such that
$T_{n+1}p_{n}=\overline{T}p_{n}$
and such that $T_{n+1}(x,\cdot)=T_{n}(x,\cdot)$ if $T_{n}p_{n}](x)=[\overline{T}p_{n}](x)$.Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing, and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the unique minimal solution of the system
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\overline{T}\overline{p}^{\mathcal{T}}$ (23)
We delay the proof of this proposition until after the following lemma.
Lemma 4.4.
For the sequence of transition matrices $(T_{n})_{n\in\mathbb{N}}$ defined in Proposition 4.3 it holds that
$\mathcal{C}_{T_{n}}\supseteq\mathcal{C}_{T_{n+1}},$
which implies that $\mathcal{C}_{T_{n}}=\mathscr{W}_{\mathcal{T}}$ for all $n\in\mathbb{N}$.
Proof.
Let $x\notin\mathcal{C}_{T_{n}}$. We will show that $x\notin\mathcal{C}_{T_{n+1}}$. If $x\in A$, then trivially $x\notin\mathcal{C}_{T_{n+1}}$, so assume $x\in A^{c}$. Then, since $x\notin\mathcal{C}_{T_{n}}$, we have $x\to^{T_{n}}y$ for some $y\in A$. By Lemma 2.14, there exists $m\in\mathbb{N}$ and a sequence of states
$x=x_{0},\ldots,x_{m}=y$
such that $T_{n}(x_{k},x_{k+1})>0$ and $p_{n}(x_{k})\leq p_{n}(x_{k+1})$ for all $k\in\{0,\ldots,m-1\}$.
If it happens that $T_{n+1}p_{n}](x_{k})=[T_{n}p_{n}](x_{k})$ for all $k\in\{0,\ldots,m\}$, then it clearly also holds that $T_{n+1}(x_{k},x_{k+1})>0$ for all $k\in\{0,\ldots,m-1\}$, and therefore $x\to^{T_{n+1}}y$ and $x\notin\mathcal{C}_{T_{n+1}}$.
On the other hand, if $T_{n+1}p_{n}](x_{k})>[T_{n}p_{n}](x_{k})$ for some $k\in\{0,\ldots,m\}$ let $k^{*}$ be the index of the first such state in the sequence $x_{0},\ldots,x_{m}$. If $x_{k^{*}}\in A$, then, similarly to the previous case, we have $x\to^{T_{n+1}}x_{k^{*}}\in A$ and we are done. Hence, we assume that $x_{k^{*}}\in A^{c}$. Next, define
$\mathcal{S}\coloneqq\{z\in\mathcal{X}:p_{n}(z)>p_{n}(x_{k^{*}})\}.$
Let us first assume ex absurdo that $\mathcal{S}=\emptyset$. Since $p_{n}(z)=1\geq p_{n}(x_{k^{*}})$ for $z\in A$, this case occurs only when $p_{n}(x_{k^{*}})=1$. In this case, we find
$1\geq[\overline{T}p_{n}](x_{k^{*}})>[T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})=1,$
where we applied our assumption that
$T_{n+1}p_{n}](x_{k^{*}})>[T_{n}p_{n}](x_{k^{*}}),$
the fact that $p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$ and our assumption that $x_{k^{*}}\in A^{c}$. Since this yields a contradiction, we have that $\mathcal{S}\neq\emptyset$.
Next, assume ex absurdo that for all $z\in\mathcal{S}$, it holds that
$T_{n+1}(x_{k^{*}},z)=0.$
Then
$[T_{n+1}p_{n}](x_{k^{*}})$ $=\sum_{z\in\mathcal{X}}T_{n+1}(x_{k^{*}},z)p_{n}(z)$
$\leq\sum_{z\notin\mathcal{S}}T_{n+1}(x_{k^{*}},z)p_{n}(x_{k^{*}})$
$=p_{n}(x_{k^{*}}),$
where we used $\sum_{z\in\mathcal{X}}T_{n+1}(x_{k^{*}},z)=1$ as $T_{n+1}$ is a transition matrix. Since $T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})$, we find $T_{n+1}p_{n}](x_{k^{*}})\leq[T_{n}p_{n}](x_{k^{*}})$ which forms a contradiction with our choice of $k^{*}$. Hence, there exists at least one $z\in\mathcal{S}$ such that $T_{n+1}(x_{k^{*}},z)>0$. Moreover, since $x\to^{T_{n+1}}x_{k^{*}}$ and $T_{n+1}(x_{k^{*}},z)>0$, we have $x\to^{T_{n+1}}z$.
So far, we have established that $x\to^{T_{n+1}}z$, with $p_{n}(x)<p_{n}(z)$. If $z\in A$, then $x\notin\mathcal{C}_{T_{n+1}}$ and we are done. Otherwise, we are back in the situation where we started, as we now want to establish that $z\to^{T_{n+1}}A$. The difference is that we have a strictly higher ‘current hitting probability’ of the state we are considering, as $p_{n}(x)<p_{n}(z)$. The state $z$ is again in one of the two cases we have already described: if it is in the first case, then $z\to^{T_{n+1}}y$ for some $y\in A$ and therefore also $x\to^{T_{n+1}}y$, which means that we can again conclude that $x\notin\mathcal{C}_{T_{n+1}}$. Otherwise, we find some $w\in\mathcal{X}$ such that $p_{n}(x)<p_{n}(z)<p_{n}(w)$ and $z\to^{T_{n+1}}w$. Since for every ‘iteration’ we do in this way we either conclude that $x\notin\mathcal{C}_{T_{n+1}}$ or strictly increase the hitting probability of the state under consideration, we must eventually be able to conclude that $x\notin\mathcal{C}_{T_{n+1}}$. Indeed, we cannot remain forever in the second case, since there are only finitely many states and each inequality in our sequence of considered hitting probabilities $p_{n}(x)<p_{n}(z)<p_{n}(w)<\dots$ is strict, meaning that we cannot repeat states. Hence we can finally conclude that $x\notin\mathcal{C}_{T_{n+1}}$, finishing the proof. ∎
Proof of Proposition 4.3.
This proof follows the structure of the proof of Proposition 4.1.
Let us first establish that the sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined. Since each $p_{n}$ is a vector of hitting probabilities, we have that $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$. Moreover, for any $n\in\mathbb{N}$, the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$
exists due to Proposition 2.5. Furthermore, for any $n\in\mathbb{N}$, the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.18.
We move on to showing that $(p_{n})_{n\in\mathbb{N}}$ is convergent and nondecreasing. We know that $p_{n}(x)=1$ for $x\in A$ for all $n\in\mathbb{N}$. Therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially nondecreasing and convergent on $A$. The same holds for $\mathscr{W}_{\mathcal{T}}$}: for all $x\in{\mathscr{W}_{\mathcal{T}}}$ and for all $n\in\mathbb{N}$ it holds that $p_{n}(x)=0$ as $p_{n}$ is the hitting probability associated with $T_{n}$ and $A$ cannot be reached with any $T\in\mathcal{T}$ if we start from $x\in{\mathscr{W}_{\mathcal{T}}}$. This means that $(p_{n})_{n\in\mathbb{N}}$ is trivially nondecreasing and convergent on $\mathscr{W}_{\mathcal{T}}$}.
We continue by showing that $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing on $A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}$. Consider any $n\in\mathbb{N}$. Note that $T_{n+1}p_{n}=\overline{T}p_{n}\geq T_{n}p_{n}$ and that Lemma 4.4 yields $\mathscr{W}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Then we have that
$p_{n}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$ $=(T_{n}p_{n})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$
$\leq(T_{n+1}p_{n})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$
$=T_{n+1}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}p_{n}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}+(T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}},$
where we used Lemma 2.6 for the last equation. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yields
$p_{n}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$ $\leq(T_{n+1}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}})^{m}p_{n}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}+\left[\sum_{k=0}^{m-1}(T_{n+1}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}})^{k}\right](T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}.$
Taking $m\to+\infty$ and applying Lemma 2.12 yields
$p_{n}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$ $\leq(I-T_{n+1})^{-1}(T_{n+1}\mathbb{1}_{A})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}$
$=p_{n+1}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}},$
where the last step follows from the fact that $p_{n+1}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n+1}p_{n+1}$. Thus, we conclude that the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing.
Since $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists and $p^{*}\in[0,1]^{{\mathcal{X}}}$. We continue by showing that $p^{*}$ satisfies the system in Equation (23). Fix any $n\in\mathbb{N}$. Then
[…]
Equation (24) now yields that
[…]
where we used Lemma 2.16(3) for the second and third inequalities. Since all summands on the final right-hand side of the equation vanish as $n$ goes to infinity, we find that $p^{*}$ solves the system in Equation (23).
Recall that $p_{n}\in[0,1]^{{\mathcal{X}}}$, $p_{n}(x)=1$ for $x\in A$ and $p_{n}(x)=0$ for $x\in\mathscr{W}_{\mathcal{T}}$ for all $n\in\mathbb{N}$. Hence also $p^{*}\in[0,1]^{{\mathcal{X}}}$, $p^{*}(x)=1$ for all $x\in A$ and $p^{*}(x)=0$ for all $x\in\mathscr{W}_{\mathcal{T}}$. Due to Proposition 3.12, we can then conclude that $p^{*}$ is the unique minimal solution to the system in Equation (19) and so $p^{*}=\overline{p}^{\mathcal{T}}$.
4.2.1 Remark - the Tiebreaking Condition
Proposition 4.3 is very similar to Proposition 4.1, but imposes an additional condition on the next transition matrix:
$T_{n+1}(x,\cdot)=T_{n}(x,\cdot)\ \text{ if }\ [T_{n}p_{n}](x)=[\overline{T}p_{n}](x).$ (25)
In other words, whenever the current row $T_{n}(x,\cdot)$ already attains the maximum $T_{n}p_{n}](x)=\max_{T\in\mathcal{T}}[Tp_{n}](x)$, the algorithm leaves that row unchanged. Equivalently, the condition acts as a tiebreaker: if several distributions achieve the maximum and $T_{n}(x,\cdot)$ is one of them, we keep $T_{n}(x,\cdot)$. This prevents pathological behaviours as illustrated in the following example.
Example 4.5.
Let $\mathcal{X}=\{1,2,3\}$ be the state space, and let $\{3\}$ be the target set. Let the set of transition matrices on $\mathcal{X}$ be
$\mathcal{T}=\operatorname{co}\left\{T=\begin{bmatrix}1&0&0\\ q_{1}&q_{2}&q_{3}\\ 0&0&1\\ \end{bmatrix}:q_{i}\in[0,1],q_{1}+q_{2}+q_{3}=1\right\}.$
Initialize the algorithm at the extreme point
$T_{1}=\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&0&1\\ \end{bmatrix}$
which has $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}=\{1\}$. This matrix maximises the hitting probability having $p_{1}=\left(0,1,1\right)^{\top}$. Any transition matrix imposing $q_{1}=0$ achieves the maximum in $\max_{T\in\mathcal{T}}[Tp_{1}](2)=q_{2}+q_{3}=1$, so one possible choice for the extreme point $T_{2}$ is
$T_{2}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}$.
However, $T_{2}$ disconnects state $2$ from the target set and its associated vector of hitting probabilities is $p_{2}=\left(0,0,1\right)^{\top}$. Therefore, without the additional tiebreaking condition in Equation (25) the algorithm might fail to converge.
5 Practical Considerations and Complexity Analysis
It remains to discuss how the algorithms described in Proposition 4.1 and Proposition 4.3 can be used in practice. This consists of three main aspects, namely how many iterations are required to reach convergence, the complexity of any such iteration, and how to find suitable starting points $T_{1}\in\mathcal{T}$. We assume that the set $\mathcal{T}$ is explicitly given, while the operators $\underline{T}$ and $\overline{T}$ are implicitly defined through Definition 2.17. This is the situation most common in practice, as the set $\mathcal{T}$ often characterises the uncertainty surrounding the true process by providing error bounds on the true transition probabilities. We mainly follow Krak’s complexity discussion [krak2021comphit].
5.1 Time per Iteration
The computation performed for each iteration in Proposition 4.1 and Proposition 4.3 can be divided into two parts. Firstly, we need to find the next element in the sequences $\left(p_{n}\right)_{n\in\mathbb{N}}$. By Proposition 2.13, this requires us to solve a linear system of equations, which can be done in at most $\mathcal{O}\left(N^{c}\right)$ time [Demmel2007]: the time required to multiply two $N\times N$ matrices. Currently, the best (published) estimate of $c$ is $c=2.371339$ [alman2025more].
Secondly, the next element of the sequences $\left(T_{n}\right)_{n\in\mathbb{N}}$ in these propositions needs to be computed. Hence, we investigate the complexity of finding some extreme point $T\in\mathcal{T}$ such that $\underline{T}f=Tf$ for some arbitrary $f\in\mathbb{R}^{\mathcal{X}}$, or, alternatively, finding some extreme point $T^{\prime}\in\mathcal{T}$ such that $\overline{T}f=T^{\prime}f$. The complexity of finding a suitable extreme point requires solving the optimisation problem from Definition 2.17 and therefore strongly depends on the definition of $\mathcal{T}$. Typically, finding such an extreme point will take at least $\mathcal{O}(N^{3})$ time [krak2021comphit]. In the remainder of the analysis, we will denote the complexity of finding an extreme point $T\in\mathcal{T}$ such that $\underline{T}f=Tf$ or $\overline{T}f=Tf$ for some arbitrary $f\in\mathbb{R}^{\mathcal{X}}$ by $\mathcal{O}(\operatorname{Extr}(\mathcal{T},\mathcal{X}))$. For both Proposition 4.1 and Proposition 4.3, the time per iteration is then $\operatorname{most}\mathcal{O}\left(N^{c}+\operatorname{Extr}(\mathcal{T},\mathcal{X})\right)$.
5.2 Starting Points
Let us discuss how to find a suitable $T_{1}\in\mathcal{T}$ for Proposition 4.1 and Proposition 4.3.
5.2.1 Lower Hitting Probabilities
As already presented in Section 4.1, definition (LR1) provides a practical algorithm to identify the set $\mathcal{A}_{\mathcal{T}}$. The computational cost of this algorithm is $\mathcal{O}\left(\operatorname{Extr}(\mathcal{T},\mathcal{X})N\right)$. This follows from the fact that for each $z\in\mathcal{X}\setminus\mathcal{D}_{k}$ we need to check whether $\underline{T}\mathbb{1}_{\mathcal{D}_{k}}](z)>0$: this operation is executed $\mathcal{O}(N)$ times.
Having obtained $\mathcal{A}_{\mathcal{T}}$, it remains to find a matrix $T_{1}\in\mathcal{T}$ such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. We observe that the matrix $T\in\mathcal{T}$ found in the last iteration of the algorithm satisfies $T\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$. In summary, identifying $\mathcal{A}_{\mathcal{T}}$ and then finding $T_{1}\in\mathcal{T}$ to use in Proposition 4.1 requires at most $\mathcal{O}\left(\operatorname{Extr}(\mathcal{T},\mathcal{X})N\right)$ time.
5.2.2 Upper Hitting Probabilities
Let us now discuss how to find a suitable $T_{1}\in\mathcal{T}$ for Proposition 4.3. The matrix $T_{1}\in\mathcal{T}$ with $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$ can be constructed row-wise. In particular, for every $x\in\mathcal{X}$, choose any $T_{1}(x,\cdot)$ in the relative interior of $\mathcal{T}_{x}$, taken with respect to the affine hull of $\mathcal{T}_{x}$. In this way, $T_{1}(x,y)>0$ for all $y\in\mathcal{X}$ such that $\overline{T}\mathbb{1}_{\{y\}}](x)>0$. The computational cost of this operation is $\mathcal{O}(N)$.
5.3 Number of Iterations
Let us now continue by investigating the number of iterations needed for convergence. We start by defining a sequence of vectors $\left(p_{n}\right)_{n\in\mathbb{N}}$ to be strictly decreasing if $p_{n+1}\leq p_{n}$ and $p_{n+1}\neq p_{n}$ for all $n\in\mathbb{N}$ and strictly increasing if $p_{n+1}\geq p_{n}$ and $p_{n+1}\neq p_{n}$. We then have the following result.
Corollary 5.1.
Let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.1.Then $(p_{n})$ is either strictly decreasing at every coordinate, or there exists $n^{*}\in\mathbb{N}$ such that $p_{n+1}<p_{n}$ for all $n<n^{*}$ and $p_{k}=p_{*}$ for all $k\geq n^{*}$, where $p_{*}=\lim_{n\to\infty}p_{n}$.
Similarly, let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.3. Then $(p_{n})$ is either strictly increasing at every coordinate, or there exists $n^{*}\in\mathbb{N}$ such that $p_{n+1}>p_{n}$ for all $n<n^{*}$ and $p_{k}=p_{*}$ for all $k\geq n^{*}$, where $p_{*}=\lim_{n\to\infty}p_{n}$.
Proof.
Completely analogous to the proof of Krak’s Corollary 6 [krak2021comphit]. ∎
If we now assume that $\mathcal{T}$ has at most $m\in\mathbb{N}$ extreme points, we can use Corollary 5.1 to find an upper bound on the number of iterations performed in Proposition 4.1 and Proposition 4.3.
Corollary 5.2.
Let $\mathcal{T}$ have finitely many extreme points; write $m\in\mathbb{N}$ for their number. Then the sequence $(p_{n})_{n\in\mathbb{N}}$ constructed in Proposition 4.1 stabilises in at most $m$ steps: there exist $n\in\mathbb{N}$ with $n\leq m$ such that $p_{k}=p_{*}$ for all $k\geq n$.
The same conclusion holds for the sequence constructed in Proposition 4.3.
Proof.
Completely analogous to the proof of Krak’s Corollary 7 [krak2021comphit]. ∎
The bound established by Corollary 5.2 is tight, i.e. there are instances in which the presented algorithms need to cycle over every extreme point of $\mathcal{T}$ before reaching the optimal transition matrix.
Example 5.3.
Let $\mathcal{X}=\{1,2,3\}$ be the state space, and let $\{2\}$ be the target set. Let the set of transition matrices be
$\mathcal{T}(m)=\operatorname{co}\left\{T_{n}=\begin{bmatrix}1-\frac{1}{n}-\frac{1}{n^{2}}&\frac{1}{n}&\frac{1}{n^{2}}\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}:n=2^{k},k\in\mathbb{N},1\leq k\leq m\right\},$
where $m\in\mathbb{N}$ is the number of extreme points of $\mathcal{T}(m)$. The situation is represented by the following graph:
[…]
Figure 1: Credal set of the first row of $\mathcal{T}(6)$.
Given $k\in\mathbb{N}$ and $n=n(k)=2^{k}$, we find the vector of hitting probabilities associated with $T_{n(k)}$ equal to $p_{n(k)}=(\frac{n}{n+1},1,0)$. We argue that the matrix $T_{n(k+1)}\in\mathcal{T}(m)$ satisfies $T_{n(k+1)}=\operatorname*{argmax}_{T\in\mathcal{T}}Tp_{n(k)}$, i.e.
$k+1=\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}T_{n(j)}p_{n(k)}.$
Expanding this expression:
$\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}T_{n(j)}p_{n(k)}$ $=\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}\left(1-2^{-j}-2^{-2j},2^{-j},2^{-2j}\right)\cdot\begin{pmatrix}\frac{2^{k}}{2^{k}+1}\\ 1\\ 0\end{pmatrix}$
$=\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}\left(\frac{2^{k}}{2^{k}+1}(1-2^{-j}-2^{-2j})+2^{-j}\right)$
$=\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}\left(2^{-j}\left(1-\frac{2^{k}}{2^{k}+1}\right)-2^{-2j}\frac{2^{k}}{2^{k}+1}\right)$
$=\operatorname*{argmax}_{j\in\mathbb{N},1\leq j\leq m}\left(2^{-j}\left(\frac{1}{2^{k}+1}-2^{-j}\frac{2^{k}}{2^{k}+1}\right)\right).$
The maximum is achieved when $2^{-j}=\frac{2^{-k}}{2}=2^{-(k+1)}$, so when $j=k+1$, which is what we wanted to show.
Initialize the algorithm for upper hitting probabilities with the matrix $S_{1}\coloneqq T_{n(1)}$. After one iteration, we get $S_{2}=T_{n(2)}$, as $S_{2}p_{n(1)}=\overline{T}p_{n(1)}$. Thus, only after cycling over all $m$ extreme points of $\mathcal{T}(m)$ we eventually find the maximising matrix in $S_{m}=T_{n(m)}$.
An analogous example can be given for lower hitting probabilities. Despite these examples, the average behaviour differs from the worst-case scenario that we just presented; the following section showcases this.
6 Numerical Analysis
In this section, we present the results of numerical experiments designed to showcase the average behaviour of the presented algorithms, namely the fact that the average number of iterations is significantly lower than the worst-case scenario illustrated in Example 5.3.
6.1 Description of the Experimental Setup
First, we explain the experimental setup. Given $N$ the number of states of $\mathcal{X}$, we set $A=\{1\}$ to be the target and construct a random graph with $N$ nodes where each directed edge indicates a positive upper probability of transition between the corresponding states. We construct the random graph in the following steps.
  • 1.
    State $N$ is made absorbing, i.e. any random walk that reaches $N$ remains there with probability one. We impose this condition so that the upper hitting probability may be less than one.
  • 2.
    We construct a random tree rooted in the target set $A$. Specifically, let $\pi$ be a random permutation of $\mathcal{X}\setminus(A\cup\{N\})$; for each state $x$ in the order given by $\pi$, we draw a directed edge from $x$ to either $A$ or to a state chosen uniformly from the set of states that have already been connected to the target set. This construction is chosen so that every state but $N$ has upper hitting probability greater than zero.
  • 3.
    We connect any two nodes of $\mathcal{X}\setminus(A\cup\{N\})$ with probability $\nicefrac{{(\lambda-1)}}{{(N-1)}}$, where $1\leq\lambda\leq N$. This Erdős-Rényi-like [erdos1960evolution] construction ensures that each node has an average degree of $\lambda$. In this way, we can easily control the graph’s density and study its influence on the average number of iterations.
We now generate the set of transition matrices $\mathcal{T}$. To do that, we propose the following two methods:
  1. 1.
    Given a parameter $0<\varepsilon<1$, we sample a transition matrix $\tilde{T}$ uniformly at random and $\varepsilon$-contaminate it [Huber1964, itip:specialcases]. In particular, the set of transition matrices is T={(1-ε)~T+εS ∣ for all trans. mat SS}. For numerical experiments, we set $\varepsilon=0.1$; experiments with other values exhibited a similar behaviour.
  2. 2.
    We sample $n_{x}$ points uniformly at random from the $n_{x}-1$-dimensional simplex, where $n_{x}$ is the number of outgoing edges of $x\in\mathcal{X}$, and $\mathcal{T}_{x}$ is their convex hull. Since the set of transition matrices $\mathcal{T}$ has separately specified rows, $\mathcal{T}$ is the Cartesian product of all $\mathcal{T}_{x}$.
6.2 Numerical Results
Since the algorithms from Propositions 4.1 and 4.3 exhibit nearly identical iteration counts, we will often present the number of iterations without explicitly specifying which algorithm it refers to. Figure 2 reports the average number of iterations, on the vertical axis, as a function of the average degree $\lambda$ of the graph, on the horizontal axis. The graphs contain either 10 (left) or 50 (right) nodes, and the set of transition matrices is generated using $\varepsilon$-contamination (top) or by taking the convex hull of random samples (bottom). The averages are obtained by running 1000 independent simulations.
Figure 2: The red and blue lines represent the average number of iterations needed to compute lower and upper hitting probabilities, respectively, as a function of the average degree of each node of the graph, $\lambda$. The two magenta shaded regions correspond to the intervals within one standard deviation ($\pm\sigma$) and approximately $95\%$ confidence ($\pm 1.96\sigma$) around the mean number of iterations.
Both algorithms are highly efficient, requiring only few iterations to converge. The average number of iterations never exceeds six, while the maximum observed was nine. The histogram in Figure 3 illustrates in detail the empirical distribution of iteration counts over the 1000 simulated experiments.
Figure 3: Number of instances, out of 1000, on which the algorithm for upper hitting probabilities converged in a given number of iterations, for different values of the average degree $\lambda$. The graph has $N=50$ nodes and the set of transition matrices is generated by taking the convex hull of transition matrices sampled uniformly at random.
Furthermore, we note that the number of iterations for the complete graph with $N$ nodes and set of transition matrices generated via $\varepsilon$-contamination is always two. This is a consequence of the fact that the credal sets of this uncertainty model are probability intervals of width $\varepsilon$, namely
$\mathcal{T}=\{T\text{ trans. mat}\mid(1-\varepsilon)\tilde{T}(x,y)\leq T(x,y)\leq(1-\varepsilon)\tilde{T}(x,y)+\varepsilon,\ \forall x,y\in\mathcal{X}\},$
for a $\tilde{T}$ sampled uniformly at random. For a node $x$ connected to the target set $A=\{1\}$, the first iteration of the algorithm that computes upper hitting probabilities assigns maximal probability of reaching $A$ in one step, i.e.
$T_{2}(x,1)=(1-\varepsilon)\tilde{T}(x,1)+\varepsilon.$
The rest of the probability mass is
$1-\left[(1-\varepsilon)\tilde{T}(x,1)+\varepsilon\right]$ $=1-\left[(1-\varepsilon)\left(1-\sum_{y\neq 1}\tilde{T}(x,y)\right)+\varepsilon\right]$
$=(1-\varepsilon)\sum_{y\neq 1}\tilde{T}(x,y)=\sum_{y\neq 1}(1-\varepsilon)\tilde{T}(x,y).$
Therefore, we must have
$T_{2}(x,y)=(1-\varepsilon)\tilde{T}(x,y),$
for all $y\in A^{c}$. The second and last iteration computes the upper hitting probability by solving the linear system and checks that no further improvement can be made. The same happens with lower hitting probabilities with $\{N\}$ in place of $\{1\}$.
Moreover, we observe a nonmonotone relationship between the density of the graph and the average number of iterations in the $\varepsilon$-contamination case. For fixed $\varepsilon$ and $N$, there is an intermediate average degree $\lambda^{*}$ that maximises the average number of iterations. This behaviour is also shown in Figure 4.
Figure 4: Average number of iterations as a function of the number of states $N$ and the average degree $\lambda$. Each red dot on the blue line marks, for a fixed number of states $N$, the average degree $\lambda^{*}$ that attains the maximum average number of iterations.
A similar behaviour is observed when probability intervals are generated uniformly at random, resulting in arbitrary widths rather than the fixed width $\varepsilon$. Although a formal theoretical justification is lacking, one plausible explanation for this empirical behaviour is the “propagation of information”. We already noted that if the graph is fully connected, the algorithm converges in two iterations: each node $x$ “knows” which of its neighbours is the one to assign maximal probability, i.e. the target or the absorbing state $N$. This does not happen in sparser graphs: not all nodes are connected, so node $x$ cannot a priori determine which neighbour should be assigned the highest probability. The following example illustrates this phenomenon.
Example 6.1.
Let $\mathcal{X}=\{1,\dots,N+2\}$ be the state space, and let $A=\{1\}$ be the target set. Let the set of transition matrices be
$\mathcal{T}=\left\{T=\begin{bmatrix}1&0&0&\dots&\dots&\dots&0\\ q_{2}&0&0&\dots&\dots&\dots&1-q_{2}\\ 0&q_{3}&0&\dots&\dots&1-q_{3}&0\\ 0&0&q_{4}&0&\dots&1-q_{4}&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\dots&\dots&0&q_{N}&1-q_{N}&0\\ \nicefrac{{1}}{{2}}&0&\dots&\dots&\dots&0&\nicefrac{{1}}{{2}}\\ 0&\dots&\dots&\dots&\dots&0&1\end{bmatrix}:\begin{aligned} &\;\;q_{x}\in[0,b],\\ &\sqrt[N]{\nicefrac{{1}}{{2}}}<b<1\end{aligned}\right\}.$
The situation is represented by the following graph:
[…]
Initialise the algorithm to compute upper hitting probabilities with any matrix $T_{1}$ satisfying $0<q_{2}<\nicefrac{{1}}{{2}}$. Then, $p_{1}(2)<\nicefrac{{1}}{{2}}$ while trivially $p_{1}(1)=1$, $p_{1}(N+2)=0$, $p_{1}(N+1)=\nicefrac{{1}}{{2}}$, and $p_{1}(x)\leq\nicefrac{{1}}{{2}}$ for every other $x\in\mathcal{X}$.
In the second iteration, node $3$ sets $q_{3}=0$ because it is connected to nodes $2$ and $N+1$, and $b\cdot p_{1}(2)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. The same argument applies for every other node $x\in\{4,\dots,N\}$: $q_{x}$ is set to zero because $b\cdot p_{1}(x-1)<p_{1}(N+1)=\nicefrac{{1}}{{2}}$. Meanwhile, node $2$ assigns probability $b$ to the transition to the target set $\{1\}$. Henceforth, solving the linear system yields $p_{2}(2)=b$ and $p_{2}(x)=\nicefrac{{1}}{{2}}$, for all $x\in\{3,\dots,N\}$.
In the third iteration, we have $b^{2}>p_{2}(N+1)=\nicefrac{{1}}{{2}}$, so node $3$ now sets $q_{3}=b$, yielding a hitting probability of $p_{3}(3)=b^{2}$. The information that larger $q_{x}$ yields greater hitting probabilities propagates backward one node per iteration, so that only after $N$ iterations the algorithm has reached convergence.
7 Conclusions and Future Work
In this work, we have investigated lower and upper hitting probabilities for imprecise Markov chains and developed efficient iterative algorithms for their computation. Here we summarise the main contributions and their implications.
In Section 3.1 we have discussed in detail the notion of reachability in the imprecise setting. The literature review highlighted several different definitions of lower reachability; hence we have explored the connections among them, presenting examples to clarify their logical relations. In particular, we have shown the equivalence between definitions $\operatorname{(LR0)}$ and $\operatorname{(LR1)}$; this equivalence has proved crucial, as the latter definition provides a practical algorithm to identify the states that cannot lower reach the target and thus have zero lower hitting probability. Using the weakest form of lower reachability, $\operatorname{(LR0)}$, we have been able to partition the state space and characterise lower hitting probabilities as the unique solution of a nonlinear fixed-point equation. An analogous characterisation has been proved for upper hitting probabilities, as they too are the unique solution to a similar fixed-point equation.
In Section 4 we have presented novel efficient iterative algorithms that compute lower and upper hitting probabilities. These algorithms alternate between solving a linear system over nontrivial states and selecting extreme points from the set of transition matrices. Although in principle these algorithms may visit all extreme points, this has never occurred in practice during numerical experiments. In fact, our experiments, described in Section 6, have shown the efficiency of the algorithms: for every graph structure and uncertainty model analysed, very few iterations were required for convergence.
We highlight several directions for future work that follow directly from our study:
  • 1.
    Formal justification of the nonmonotone relation between density and number of iterations. We would like to investigate why the algorithms for computing lower and upper hitting probabilities show a nonmonotone relation between the density $\lambda$ of the graph and the average number of iterations. For specific uncertainty models (e.g. $\varepsilon$-contamination, probability intervals, etc.), we may try to find the average degree $\lambda^{*}$ that maximises the number of iterations.
  • 2.
    Handling pathological credal sets. Although extremely rare, some credal set shapes force the algorithms to cycle over all extreme points, as in Example 5.3. Future work should identify such cases in advance and aim to derive specialised algorithms to compute lower and upper hitting probabilities, that perform better in such regimes.
  • 3.
    Extension to continuous-time IMCs. One relatively straightforward step forward, would be to generalise the fixed-point characterisation and the computational framework discussed here, to continuous-time imprecise Markov chains.
  • 4.
    Lower and upper meeting probabilities. It would be interesting to characterise and compute efficiently lower and upper meeting probabilities, i.e. the probability that two (or more) processes are ever in the same state at the same time. This present paper, combined with Sangalli et al.’s recent work on meeting times for imprecise Markov chains [sangalli2025meeting], provides a foundation for this line of research.
  • 5.
    Relation between lower and upper hitting times and probabilities. Finally, we may aim to investigate the relation between hitting times and probabilities in the imprecise framework, and seek to establish whether knowledge of one quantity allows us to infer bounds on the other.
Acknowledgements
This work has been partly supported by the PersOn project (P21-03), which has received funding from Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO).
Appendix A Proofs of Auxiliary Lemmas
Proof of 2.6.
Consider any $x\in A^{c}\setminus\mathcal{S}$. Then
$[Tf](x)$ $=\sum_{y\in\mathcal{X}}T(x,y)f(y)$
$=\sum_{y\in\mathcal{X}}T(x,y)\Bigl[\bigl(f|_{A^{c}\setminus S}\bigr)^{\uparrow_{\mathcal{X}}}(y)+\mathbb{1}_{A}(y)\Bigr]$
$=\Bigl[T\Bigl(\bigl(f|_{A^{c}\setminus\mathcal{S}}\bigr)^{\uparrow_{\mathcal{X}}}\Bigr)\Bigr](x)+[T\mathbb{1}_{A}](x)$
$=\bigl[T|_{A^{c}\setminus\mathcal{S}}\,(f|_{A^{c}\setminus\mathcal{S}})\bigr](x)+[T\mathbb{1}_{A}](x),$
where we used $f(y)=1$ for $y\in A$ and $f(y)=0$ for $y\in\mathcal{S}$. ∎
Proof of Lemma 2.7.
Observe that, by definition
$T^{n}\mathbb{1}_{A}](x)=\sum_{z\in\mathcal{X}}T(x,z)\,[T^{\,n-1}\mathbb{1}_{A}](z).$
All terms in the sum are nonnegative and by hypothesis the left-hand side is strictly greater than 0. Hence there exists $y\in\mathcal{X}$ such that
$T(x,y)\,[T^{\,n-1}\mathbb{1}_{A}](y)>0,$
so in particular $T(x,y)>0$ and $T^{\,n-1}\mathbb{1}_{A}](y)>0$.
We now check that $y\in A^{c}\setminus\mathcal{C}_{T}$. First, $y\notin A$: if $y\in A$ then $T(x,y)>0$ would imply $T\mathbb{1}_{A}](x)\geq T(x,y)>0$, contradicting the minimality assumption that $T^{m}\mathbb{1}_{A}](x)=0$ for every $m<n$ (recall $n>1$). Thus $y\in A^{c}$.
Second, $y\notin\mathcal{C}_{T}$: by definition any $z\in\mathcal{C}_{T}$ satisfies $T^{m}\mathbb{1}_{A}](z)=0$ for all $m\in\mathbb{N}$, but we have $T^{\,n-1}\mathbb{1}_{A}](y)>0$, so $y\notin\mathcal{C}_{T}$.
Therefore $y\in A^{c}\setminus\mathcal{C}_{T}$, $T(x,y)>0$ and $T^{\,n-1}\mathbb{1}_{A}](y)>0$, which proves the lemma. ∎
Proof of Lemma 2.8.
We first prove the monotonicity claim. Let $M\coloneqq T|_{\mathcal{S}}$. If $f\leq g$ then for every $x\in\mathcal{S}$
$Mf](x)=\sum_{y\in\mathcal{S}}M(x,y)f(y)\leq\sum_{y\in\mathcal{S}}M(x,y)g(y)=[Mg](x),$
since each $M(x,y)\geq 0$. Thus $Mf\leq Mg$. Iterating this argument gives, by simple induction on $k$, that $M^{k}f\leq M^{k}g$ for all $k\in\mathbb{N}_{0}$.
We now prove the bounds for $\mathbf{0}$ and $\mathbf{1}$. First note that $\mathbf{0}\leq\mathbf{1}$, so by the monotonicity just proved we have
[…]
But $M^{k}\mathbf{0}=\mathbf{0}$ for every $k$. Hence $\mathbf{0}\leq M^{k}\mathbf{1}$.
For the upper bound, observe that for every $x\in\mathcal{S}$
$M\mathbf{1}](x)=\sum_{y\in\mathcal{S}}M(x,y)\leq\sum_{y\in\mathcal{X}}T(x,y)=1.$
Thus $M\mathbf{1}\leq\mathbf{1}$. Applying the monotonicity repeatedly yields
$M^{k}\mathbf{1}=M^{k-1}(M\mathbf{1})\leq M^{k-1}\mathbf{1}\leq\cdots\leq M\mathbf{1}\leq\mathbf{1},$
so $M^{k}\mathbf{1}\leq\mathbf{1}$ for every $k\in\mathbb{N}_{0}$.
Combining the two inequalities gives $\mathbf{0}\leq M^{k}\mathbf{1}\leq\mathbf{1}$ for all $k\in\mathbb{N}_{0}$, which completes the proof. ∎
Proof of Lemma 2.9.
Let $M\coloneqq T|_{{A^{c}\setminus\mathcal{C}_{T}}}$ and fix $x\in{A^{c}\setminus\mathcal{C}_{T}}$. Since $x\notin\mathcal{C}_{T}$, there exists a (finite) minimal integer
$n_{x}\coloneqq\min\{n\geq 1:\,[T^{n}\mathbb{1}_{A}](x)>0\}.$
Consider any path of length $n_{x}$, $x_{0},\dots,x_{n_{x}}$, with $x_{0}=x$ and $x_{n_{x}}\in A$. Because $m<n_{x}$ implies $T^{m}\mathbb{1}_{A}](x)=0$, such a path cannot visit $A$ before time $n_{x}$. Moreover, by the definition of $\mathcal{C}_{T}$, a path that ever enters $\mathcal{C}_{T}$ cannot reach $A$ at any later time; hence every path contributing to $T^{n_{x}}\mathbb{1}_{A}](x)$ stays in $A^{c}\setminus\mathcal{C}_{T}$} up to time $n_{x}-1$ and then jumps from $A^{c}\setminus\mathcal{C}_{T}$} to $A$ at time $n_{x}$.
On the other hand, $M^{n_{x}}{\mathbf{1}}](x)$ is the total probability of all length-$n_{x}$ paths that remain in $A^{c}\setminus\mathcal{C}_{T}$} for all $n_{x}$ steps. Therefore the two events are disjoint, and we obtain
$M^{n_{x}}{\mathbf{1}}](x)\leq 1-[T^{n_{x}}\mathbb{1}_{A}](x)<1.$
Thus, for every $x\in{A^{c}\setminus\mathcal{C}_{T}}$, there exists $n_{x}\geq 1$ such that $M^{n_{x}}{\mathbf{1}}](x)<1$.
Now we use Lemma 2.8: since $M{\mathbf{1}}\leq{\mathbf{1}}$, the map $k\mapsto M^{k}{\mathbf{1}}$ is componentwise nonincreasing: $M^{k+1}{\mathbf{1}}=M^{k}(M{\mathbf{1}})\leq M^{k}{\mathbf{1}}$ for all $k\in\mathbb{N}_{0}$. Let $n\coloneqq\max_{x\in{A^{c}\setminus\mathcal{C}_{T}}}n_{x}$, which is finite because $A^{c}\setminus\mathcal{C}_{T}$} is finite. Then for any $k\geq n$ and any $x\in{A^{c}\setminus\mathcal{C}_{T}}$,
$M^{k}{\mathbf{1}}](x)\leq[M^{n_{x}}{\mathbf{1}}](x)<1.$
This proves the claim. ∎
Proof of Corollary 2.10.
We show by induction that $(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)\leq[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)$ for all $x\in A^{c}\setminus\mathcal{S}$ and all $k\in\mathbb{N}$.
First consider $k=1$. Then
$[T|_{A^{c}\setminus\mathcal{S}}\mathbf{1}](x)$ $=[T\mathbb{1}_{A^{c}\setminus\mathcal{S}}](x)$
$\leq[T\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T}}](x)$
$=[T|_{A^{c}\setminus\mathcal{C}_{T}}\mathbf{1}](x),$
where the inequality follows from the linearity of $T$ and the fact that we assumed $\mathcal{C}_{T}\subseteq\mathcal{S}$.
Now let us assume as induction hypothesis that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}\Bigr](y)\leq\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}\Bigr](y)$
for all $y\in A^{c}\setminus\mathcal{S}$ for some $k\in\mathbb{N}$. We will show that the same statement then also holds for $k+1$. We find that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k+1}\mathbf{1}\Bigr](x)$ $=[T|_{A^{c}\setminus\mathcal{S}}(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}](x)$
$=\Bigl[T\Bigl(((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](x)$
$=\sum_{y\in\mathcal{X}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\sum_{y\in A^{c}\setminus\mathcal{S}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$\leq\sum_{y\in A^{c}\setminus\mathcal{S}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y),$
where the inequality follows from the induction hypothesis. We now add the terms of $(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}$ corresponding to the states in $A^{c}\cap(\mathcal{S}\setminus\mathcal{C}_{T})$ to the sum, which are nonnegative due to Lemma 2.8. This then yields that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k+1}\mathbf{1}\Bigr](x)$ $\leq\sum_{y\in A^{c}\setminus\mathcal{C}_{T}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\sum_{y\in\mathcal{X}}T(x,y)\Bigl[((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](y)$
$=\Bigl[T\Bigl(((T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1})^{\uparrow\mathcal{X}}\Bigr](x)$
$=[T|_{A^{c}\setminus\mathcal{C}_{T}}\,(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}](x)$
$=\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k+1}\mathbf{1}\Bigr](x).$
This completes the induction proof. Hence we have shown that
$\Bigl[(T|_{A^{c}\setminus\mathcal{S}})^{k}\mathbf{1}\Bigr](x)\leq\Bigl[(T|_{A^{c}\setminus\mathcal{C}_{T}})^{k}\mathbf{1}\Bigr](x)$
for all $x\in A^{c}\setminus\mathcal{S}$ and all $k\in\mathbb{N}$. Applying Lemma 2.9 now yields the result. ∎
Proof of Corollary 2.11.
Let $\mathcal{S}\subseteq\mathcal{X}$ satisfy $\mathcal{C}_{T}\subset\mathcal{S}\subset A^{c}$ and set
$M\coloneqq T|_{A^{c}\setminus\mathcal{S}}.$
By Corollary 2.10 there exists $n\in\mathbb{N}$ such that
$M^{n}\mathbf{1}](x)<1\quad\text{for all }x\in A^{c}\setminus\mathcal{S}.$
Then, Lemma 2.8 yields $0\leq M^{n}\mathbf{1}$. Thus
$M^{n}\mathbf{1}||<1.$
[…]
$M^{n}\mathbf{1}=M^{n}(-\mathbf{1})\leq M^{n}f\leq M^{n}\mathbf{1}.$
[…]
which proves the claim. ∎
Proof of Lemma 2.14.
Since $x\rightarrow^{T}y$ for some $y\in A$, we have $c\coloneqq p^{T}(x)>0$. Suppose $c=1$. Then, either $x\in A$ and we are done, or $x$ is surrounded by states with hitting probability 1, i.e. $p^{T}(w)=1$ for all $w\in\mathcal{X}$ such that $T(x,w)>0$. From the finiteness of $\mathcal{X}$, it follows that there exists the desired path.
If $0<c<1$, then define the set of states reachable from $x$ with hitting probability equal to $c$:
$C\coloneqq\{z\in\mathcal{X}:x\rightarrow^{T}z,\ p^{T}(z)=c\}.$
This set is not closed: if it were the case, every state in $C$ could not reach $A$ and would have hitting probability equal to zero. Using the fact that
$p^{T}(z)=\sum_{w\in\mathcal{X}}T(z,w)p^{T}(z),$
we obtain that either some neighbour $v\in\mathcal{X}$ of $z\in C$ has $p^{T}(v)>c$ or all neighbours of $z$ are in $C$. Therefore, there exist $m\in\mathbb{N}$ and a sequence $x=x_{0},\dots,x_{m}$ such that $T(x_{i},x_{i+1})>0$ and $p^{T}(x_{i})=c<p^{T}(x_{m})$, for all $i=0,\dots,m-1$. Since $\mathcal{X}$ is finite, repeating this argument at most $N={\mathcal{X}}$ times gets the desired sequence of states. ∎